### v1

# Daily-Omni: Towards Audio-Visual Reasoning with Temporal Alignment across Modalities

Abstract
Recent Multimodal Large Language Models (MLLMs) achieve promising performance on visual and audio benchmarks independently. However, the ability of these models to process cross-modal information synchronously remains largely unexplored. In this paper, we introduce: 1) Daily-Omni, an Audio-Visual Questioning and Answering benchmark comprising 684 videos of daily life scenarios from diverse sources, rich in both audio and visual information, and featuring 1197 multiple-choice QA pairs across 6 major tasks; 2) Daily-Omni QA Generation Pipeline, which includes automatic annotation, QA generation and QA optimization, significantly improves efficiency for human evaluation and scalability of the benchmark; 3) Daily-Omni-Agent, a training-free agent utilizing open-source Visual Language Model (VLM), Audio Language Model (ALM) and Automatic Speech Recognition (ASR) model to establish a baseline for this benchmark. The results show that current MLLMs still struggle significantly with tasks requiring audio-visual integration, but combining VLMs and ALMs with simple temporal alignment techniques can achieve substantially better performance. Codes and benchmark are available at https://github.com/Lliar-liar/Daily-Omni.
1 Introduction
Non-textual modalities, such as vision and hearing, often convey richer and more nuanced information than text alone in daily life scenarios, playing a crucial role in our understanding of and interaction with the physical world. Therefore, advancements in Multimodal Large Language Models (MLLMs) capable of comprehensively understanding the multi-modal information are a crucial foundation for achieving artificial intelligence capable of interacting with the physical world and observing the outcome of its operations.
Recent MLLMs [xu_qwen25-omni_2025, sun_video-salmonn_2024, noauthor_introducing_2024, zhang_internlm-xcomposer25-omnilive_2024, fu_vita-15_2025, liu_ola_2025, cheng_videollama_2024, li_baichuan-omni-15_2025, xie_mini-omni2_2024, guo_m2-omni_2025, team_reka_2024, microsoft2025phi4minitechnicalreportcompact, lu2023unifiedio2scalingautoregressive, liu_nexus-o_2025, gemmateam2025gemma3technicalreport] have demonstrated groundbreaking capabilities spanning audio (ASR, sound classification, captioning) and visual domains (OCR, VQA, video grounding), with significant accuracy improvements over previous benchmarks. However, existing methods still face several limitations. Firstly, many MLLMs predominantly focus on visual abilities, often neglecting the importance of other modalities like audio. This oversight may stem from the fact that current visual datasets are more abundant, of higher quality, and cover a broader range of tasks compared to audio datasets. Existing audio datasets [Librispeech, wang2020covost2massivelymultilingual, poria2019meldmultimodalmultipartydataset, chen_vggsound_2020, Audioset, Vocalsound] tend to prioritize speech-related or music-related tasks and basic sound classification, often overlooking more complex yet crucial tasks such as reasoning over generic sounds. Consequently, many MLLMs incorporate only speech encoders as their primary auditory component or rely heavily on speech-related datasets for audio pre-training. This architectural limitation fundamentally restricts their ability to comprehend rich acoustic environments where non-speech sounds (e.g., environmental noises, mechanical failures, or emotional cues in non-verbal vocalizations) carry critical semantic information. Secondly, the current landscape lacks high-quality multimodal datasets that integrate temporally aligned auditory and visual information. Existing audio-visual datasets and benchmarks [yun_pano-avqa_2021, li_learning_2022, avqa, li_omnibench_2024, hong_worldsense_2025, gong_av-odyssey_2024, sungbin2025avhbenchcrossmodalhallucinationbenchmark, geng2025longvalevisionaudiolanguageeventbenchmarktimeaware] reveal three persistent limitations. First, several focus on specialized scenarios [yun_pano-avqa_2021, li_learning_2022] such as musical performances or panoramic environments, thereby introducing domain-specific biases. Second, many employ static image-audio pairs [li_omnibench_2024, gong_av-odyssey_2024] that disregard crucial temporal dynamics inherent in real-world video contexts. Third, current tasks are often too narrow, with many benchmarks focusing on specific applications such as captioning or open-ended responses [geng2025longvalevisionaudiolanguageeventbenchmarktimeaware]. The common absence of rigorous evaluation frameworks and standardized metrics makes it hard to reliably compare results. While WorldSense [hong_worldsense_2025] established a valuable audio-visual multi-choice QA benchmark for daily scenarios through its meticulous dataset curation, two critical limitations persist: (1) it lacks a systematic framework for scalable QA generation, relying instead on manual annotation processes that hinder dataset expansion; and (2) the benchmark serves primarily as an evaluation tool, offering limited methodological guidance on enhancing model capabilities via explicit training protocols or architectural modifications. This dual limitation constrains both the benchmark’s adaptability to emerging domains and its practical utility in driving model improvements.
Figure 1: Examples of Daily-Omni QAs. The audio and visual information required for answering the questions are provided in the figure. The correct answer for the given questions are highlighted.
In this paper, we introduce Daily-Omni, an Audio-Visual Questioning and Answering benchmark with 684 videos featuring daily life scenarios from various sources [Audioset, fu2024videommefirstevercomprehensiveevaluation, Farré2024FineVideo] with rich audio and visual information and 1197 multiple choice QA-pairs across 6 major tasks ranging from audio visual event aligning to complicated cross-modal reasoning. The videos are sampled from different datasets and segmented into 30-second or 60-second intervals to systematically evaluate model performance across different temporal contexts. We further introduce Daily-Omni QA Generation Pipeline which encompasses five automated modules: video annotation, annotation revision, audio-visual temporal alignment, QA generation, and QA optimization. This framework demonstrates remarkable scalability: a single annotator can complete quality filtering process for these 1197 high-quality QA pairs within 30 hours, achieving an approximate 30% acceptance rate from initially generated candidates. In addition, we propose Daily-Omni Agent, an audio-visual agent utilizing open-source Visual Language Model (VLM), Audio Language Model (ALM) and Automatic Speech Recognition (ASR) model without further finetuning to establish a baseline for this benchmark. We evaluated recent MLLMs and our agent on the Daily-Omni benchmark, where the Daily-Omni Agent achieved state-of-the-art performance among open-source methods. The experimental results show that current MLLMs still face significant challenges in tasks requiring deep audio-visual temporal integration. They also reveal that by combining existing visual and audio language models with simple temporal alignment techniques, as demonstrated by our Daily-Omni Agent, substantially improved performance can be achieved, underscoring a promising direction for enhancing multimodal reasoning.
2 Related Works
2.1 Multimodal Large Language Models
Recent advancements in Multimodal Large Language Models (MLLMs) primarily fall into three categories: Audio Language Models (ALMs) incorporating audio capabilities [tang_salmonn_2024, gong2024listenthinkunderstand, chu_qwen-audio_2023, chu2024qwen2audiotechnicalreport, ghosh2025audioflamingo2audiolanguage, ghosh2024gamalargeaudiolanguagemodel], Visual Language Models (VLMs) adding visual understanding [liu2025nvilaefficientfrontiervisual, bai_qwen25-vl_2025, lu2024deepseekvlrealworldvisionlanguageunderstanding, wang2024emu3nexttokenpredictionneed], and Omni-modal Language Models (OLMs) combining both audio and visual modalities [noauthor_introducing_2024, li_baichuan-omni-15_2025, zhang_internlm-xcomposer25-omnilive_2024, team_reka_2024, cheng_videollama_2024, liu_ola_2025, microsoft2025phi4minitechnicalreportcompact, fu_vita-15_2025, guo_m2-omni_2025, xu_qwen25-omni_2025, liu_nexus-o_2025, gemmateam2025gemma3technicalreport]. These MLLMs often employ a modular architecture, utilizing separate encoders for audio and visual inputs. In the audio domain, encoders such as [radford2022robustspeechrecognitionlargescale, chen_beats_2022, li2024mertacousticmusicunderstanding] are commonly used to extract sound representations. Some models, such as [tang_salmonn_2024, liu_ola_2025, zhang_internlm-xcomposer25-omnilive_2024], even integrate multiple audio encoders specialized for different sound types (e.g., speech, music). Similarly, visual encoders [liu2021swintransformerhierarchicalvision, li2023blip2bootstrappinglanguageimagepretraining, dai2023instructblipgeneralpurposevisionlanguagemodels, liu2023visualinstructiontuning, dehghani2023patchnpacknavit] are used to process images and videos. However, this modular approach, while cost-efficient and effective for individual modalities, struggles to capture the crucial temporal correlations inherent in synchronized audio-visual streams. This limitation arises because audio and visual inputs are encoded independently. Although multi-modal positional embedding techniques like M-RoPE [bai2023qwenvlversatilevisionlanguagemodel] and TMRoPE [xu_qwen25-omni_2025] enhance cross-modal temporal understanding to some extent, effective methods for deeply integrating and temporally aligning multimodal data remain relatively scarce. Additionally, in several MLLMs [microsoft2025phi4minitechnicalreportcompact, fu_vita-15_2025, zhang_internlm-xcomposer25-omnilive_2024] , the audio encoder primarily serves to process user instructions–akin to how the text modality is used–rather than sensing the ambient environment, which is a common function of the visual modality.
2.2 Audio-Visual Understanding Datasets and Benchmarks
The development of uni-modal datasets and associated tasks for audio and vision has driven advancements in Audio Language Models (ALMs) and Visual Language Models (VLMs). Visual datasets and benchmarks [li2024llavanextinterleavetacklingmultiimagevideo, fu2024videommefirstevercomprehensiveevaluation, wu_star_2024, mangalam_egoschema_2023, liu2023visualinstructiontuning, wu_longvideobench_nodate, zhang2024videoinstructiontuningsynthetic, yue2024mmmumassivemultidisciplinemultimodal, yue2024mmmuprorobustmultidisciplinemultimodal, li2023seedbenchbenchmarkingmultimodalllms, fu2024ocrbenchv2improvedbenchmark, gao2017talltemporalactivitylocalization, hu2025videommmuevaluatingknowledgeacquisition] primarily focus on tasks for static images (such as OCR, grounding, segmentation, classification, and question-answering) and dynamic videos (such as captioning, temporal grounding, and understanding), while audio datasets and benchmarks [ghosh2025audioflamingo2audiolanguage, Audioset, chen_vggsound_2020, Librispeech, yang2024airbenchbenchmarkinglargeaudiolanguage] address speech-related tasks (including ASR, emotion recognition, and entity recognition) and non-speech tasks (such as sound classification and audio grounding). However, while attempts at creating audio-visual datasets date back to at least 2021 [yun_pano-avqa_2021], these early efforts often had significant limitations. For example, Music-AVQA [li2022learninganswerquestionsdynamic] focused specifically on music performance videos, while Pano-AVQA [yun_pano-avqa_2021] centered on panoramic videos. Others, such as AVQA [avqa] and OmniBench [li_omnibench_2024], were restricted to short, simple videos or static images. Furthermore, AV-Odyssey [gong_av-odyssey_2024] heavily emphasized specific audio tasks, such as recognizing timbre and loudness, rather than broader audio-visual understanding. While WorldSense [hong_worldsense_2025] provides a valuable model for real-world audio-visual question-answering, how to efficiently scale AVQA datasets and enhance OLM abilities is an area requiring further exploration.
3 Daily-Omni
This section details the Daily-Omni framework, which comprises three core components: multi-modal data curation, hierarchical video annotation, and QA synthesis and evaluation. Furthermore, we delineate the evaluation paradigm implemented through the Daily-Omni Agent, which serves as our benchmarking baseline for systematic assessment. Daily-Omni aims to establish a benchmarking framework designed to systematically evaluate MLLMs’ ability to perceive and reason in real-life audio-visual scenarios. As illustrated in Figure LABEL:Fig:qa_count, the Daily-Omni benchmark comprises videos from all 11 YouTube categories. The QAs are meticulously crafted to assess a spectrum of multi-modal capabilities, spanning from basic cross-modal perception to intricate reasoning. All questions are designed to require the integration of audio, visual, and textual information for a correct answer. In total, the Daily-Omni benchmark comprises 684 videos and 1197 QA pairs. Of these QAs, 550 correspond to 60-second videos and 647 correspond to 30-second videos. We compared the features of Daily-Omni with various audio-visual benchmarks in Table 1.
3.1 Data Curation
Daily-life scenarios offer abundant information across both audio and visual modalities. Within the visual modality, our data curation focuses on selecting videos characterized by significant temporal dynamics. Conversely, static scenes, exemplified by a vlogger addressing the camera with limited motion, offer minimal temporal information for questioning and are usually excluded. To prevent data bias, videos are selected from a wide range of categories and topics. Videos with harmful or discriminative content are excluded. Regarding the audio modality, recognizing that prior datasets have concentrated heavily on speech, our benchmark is designed to encompass a broader range of everyday sounds, such as music, speech, and other sound events. These acoustic signals should be present within the videos, appearing either concurrently or consecutively. Furthermore, to maintain focus on audio-visual processing and avoid challenges related to multilingual text comprehension, videos with languages other than English are also excluded.
Our video data originates from AudioSet [Audioset], Video-MME [fu2024videommefirstevercomprehensiveevaluation], and FineVideo [Farré2024FineVideo]. While AudioSet provides 10-second clips primarily for audio classification, we retrieved the original, full-length source videos. These were then processed into longer 30s or 60s segments which contains the original 10s segments to ensure they contain certain types of sound events. To ensure inclusion of spoken content, we employed Whisper-V3-Large [radford2022robustspeechrecognitionlargescale] to detect speech within these segments, adding those segments identified as containing speech to our set of video candidates. For Video-MME and FineVideo, our selection criteria prioritized videos containing substantial audio information alongside rich temporal visual dynamics. The inclusion of these datasets also increases diversity, as they feature more recent videos compared to Audio Set (primarily pre-2017 uploads), thus broadening the representation of genres and styles.
3.2 Data Annotation & QA Construction
We developed a pipeline that employs MLLMs to generate and revise visual and audio annotations. Concurrently, Reasoning Large Language Models (LLMs) are utilized to construct and optimize the associated questions, choices, and answers. To obtain detailed annotations while ensuring cost-effectiveness, we specifically used Gemini 2.0 Flash [noauthor_introducing_2024] for the annotation task and Deepseek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability] for QA construction and optimization. Figure 3 provides a outline of this process.
Figure 3: The outline of Daily-Omni QA construction pipeline. The arrows indicates the sequence of the processes.
Segment Annotation Recognizing that even state-of-the-art Multimodal Large Language Models (MLLMs) can exhibit cross-modal hallucinations [sungbin2025avhbenchcrossmodalhallucinationbenchmark], we employed Gemini 2.0 Flash to annotate the audio and visual modalities independently. Additionally, since MLLM performance is known to degrade when processing long audio clips, we segmented the videos prior to annotation. Specifically, each clip was divided into three equal, shorter segments (e.g., 10s segments for 30s clips, 20s segments for 60s clips) to improve the MLLM’s subsequent audio annotation quality.
Visual & Audio Revision After obtaining visual annotations from the video segments (processed without audio) and audio annotations from the corresponding audio segments, we use Gemini 2.0 Flash to perform a consistency check on the visual annotations. For this step, we prompt the model with the complete video clip, allowing it to review the segment annotations and ensure overall coherence. For example, referencing the full-length video, the model verifies whether a person described in an early segment is the same individual appearing in subsequent segments and generates a consistent revised annotation. Subsequently, audio annotations undergo refinement using a Reasoning LLM (Gemini 2.0 was employed in the initial phase of the project, with Deepseek-R1 used in later stages). This model leverages the consistent visual annotations to perform cross-modal correction, rectifying sound misidentifications and identifying sound sources. For example, if the audio is annotated as a ’generic impact sound’ but the visual annotation for the same segment shows a ’door slamming shut’, the Reasoning LLM uses this visual context to correct the audio description to ’door slamming sound’ and attributes the sound’s origin to the observed door.
Visual & Audio Event Alignment At this stage, we have sequences of visual and audio events annotated within consecutive 10s or 20s segments. However, these segment-level annotations do not explicitly specify the temporal alignment between individual visual and audio events – i.e., which specific events occurred simultaneously. To establish this precise cross-modal event concurrency, we proposed a novel event aligning technique. By prompting Gemini 2.0 Flash with the complete audio-visual clip, we instruct it to identify the visual event(s) occurring concurrently with each identified audio event. These aligned audio-visual event pairs provide sufficient information to infer the temporal relationships between any audio and visual event within the sequence. The details of annotation generation, revision and event alignment is shown in Figure 4.
Figure 4: Details of Daily-Omni annotation generation, revision and event alignment. For cost-efficiency, we align all events with one query.
QA Construction Using the consistent visual annotations, revised audio annotations, and aligned event pairs derived from each video, we prompted Deepseek-R1 to generate multi-choice questions covering the following types: (1) AV Event Alignment: Questions to determine which audio and visual events occurred simultaneously with each other; (2) Event Sequence: Questions to determine the temporal sequence of visual and audio events in the video; (3) Reasoning: Questions to explain the cause or reason behind the occurrence of a visual or audio event in the video; (4) Inference: Questions to speculate on information not explicitly presented in the video; (5) Comparative: Questions to compare the similarity or difference between the audio and visual information of two or more events in the video; (6) Context Understanding: Questions to determine the contextual information surrounding a specific event in the video. Since Daily-Omni aims to evaluate MLLM perception and reasoning within real-life audio-visual scenarios, we deliberately exclude certain question types. Specifically, counting and measuring questions are omitted, as they often rely on a single modality rather than integrated multi-modal understanding. Similarly, purely knowledge-based queries, such as celebrity identification, fall outside our intended scope.
Table 1: Comparison of various audio-visual benchmarks. The table details key characteristics including modality (A: audio, V: video, I: image), dataset size (#Video, #QA Pairs), and Question Type (MC: multiple choice, DF: defined word, BB: bounding boxes). ’Efficient Scalability’ indicates whether the associated work proposed an automated method for benchmark expansion. ’Open-Domain’ denotes benchmarks featuring various categories visual content, while ’General Sound’ signifies benchmarks whose audio content includes speech, music, and generic real-life sounds.
Benchmarks Modality #Video #QA Pairs Question Type Efficient Scalability Open-Domain General Sound
AVQA [avqa] V+A 57,015 57,335 MC
Music-AVQA [li2022learninganswerquestionsdynamic] V+A 9,288 45,867 DF
Pano-AVQA [yun_pano-avqa_2021] V+A 5,400 51,700 DF & BB
OmniBench [li_omnibench_2024] I+A 1,142 MC
AV-Odyssey [gong_av-odyssey_2024] I+A 4,555 MC
WorldSense [hong_worldsense_2025] V+A 1,662 3,172 MC
Daily-Omni (ours) V+A 684 1197 MC
QA Optimization and Quality Control By avoiding strict question templates, the generated QAs exhibit greater creativity, incorporate complicated logic, and sometimes feature obscure descriptions, making them more closely resemble questions asked in real-life scenarios. However, the generated questions and choices can sometimes contain excessive textual information, potentially allowing powerful models to infer the correct answer using text alone, without engaging with the audio-visual content. Therefore, Deepseek-R1 was employed again to remove superfluous textual information from the questions and choices. Additionally, it replaced obviously incorrect options with more challenging distractors, thereby increasing the difficulty and reducing the potential for correct answers based on guessing alone. Subsequently, we evaluated the optimized questions using two powerful LLMs, GPT-4o [openai_gpt-4o_2024] and Deepseek-V3 [deepseekai2025deepseekv3technicalreport], providing them with only the textual questions and choices (no audio-visual context). Questions that could be answered correctly by both LLMs under this text-only condition were discarded, as they did not necessitate multimodal reasoning. This automated filtering step resulted in approximately 47% of the candidate QAs being discarded. Finally, the remaining QAs underwent manual evaluation for quality control. Human evaluators examined each QA, verifying: (1) that there was exactly one unambiguously correct answer among the choices, (2) that the proposed answer was indeed the correct one, and (3) that answering the question genuinely required comprehensive audio-visual capabilities. Based on this assessment, evaluators either accepted the QA for inclusion in the final benchmark or rejected it. Facilitated by the automated pipeline, the final human evaluation process was highly efficient. A single annotator used less than 30 hours to review the candidates and establish the final set of 1197 QAs, corresponding to an acceptance rate of approximately 30% during this manual review stage.
Figure 5: The outline of Daily-Omni Agent.
3.3 Daily-Omni Agent
To establish a baseline for MLLMs and demonstrate the importance of temporal awareness in audio-visual question answering, we constructed the "Daily-Omni Agent". This agent, designed to understand audio-visual context and answer related questions, integrates several models: Qwen2-Audio (7B parameters) [chu2024qwen2audiotechnicalreport], Qwen2.5-VL-7B [bai_qwen25-vl_2025], Whisper-Large-V2 [radford2022robustspeechrecognitionlargescale], and the text-based Qwen2.5-14B-Instruct [qwen2025qwen25technicalreport]. As shown in Figure 5, when presented with a question, its choices, and the associated video context, the agent first divides both the video and audio streams into three segments of equal duration. Subsequently, it independently generates annotations for these segments: visual annotations using Qwen2.5-VL and audio annotations using Qwen2-Audio. Additionally, we use Whisper-Large-V2 to provide a transciption of speech and singing of each segment, as Qwen2-Audio tends to omit this information in its annotations. Following the segment annotation, the agent utilizes Qwen2.5-VL to perform a consistency check on the visual annotations using the complete video, mirroring the revision process described previously (Section 3.2).
While generating aligned audio-visual event pairs would ideally provide richer temporal information for reasoning, implementing this step within the agent presents practical challenges. Unlike the data curation pipeline, the agent lacks access to a highly capable OLM like Gemini 2.0 Flash for precise event alignment. Furthermore, providing a large number of fine-grained event pairs risks overwhelming the context capacity or reasoning capabilities of the agent’s LLM. Therefore, as an alternative to full event alignment, we adopt a innovative targeted approach: First, we prompt Qwen2.5-14B-Instruct with the visual and audio annotations, the question, and its choices. The model’s task is to identify a list of specific visual events whose temporal localization is deemed necessary to answer the question correctly. Subsequently, we utilize Qwen2.5-VL-7B, functioning as a video temporal grounding model, to determine the start and end timestamps for each of these identified critical events. If the duration of a grounded event interval falls below a predefined threshold, the agent classifies this interval as critical. It then retrieves descriptions of both visual and audio events occurring within this specific, brief period using the previously mentioned approach, thereby establishing a localized alignment between concurrent events. Finally, we prompted Qwen2.5-14B-Instruct to determine the correct answer by providing it with the question, choices, the previously generated visual and audio annotations, and the extracted aligned event pairs.
4 Experiment
In this section, we present a comprehensive evaluation of recent open-source and proprietary Multimodal Large Language Models (MLLMs) on our benchmark to delineate their capability boundaries. Furthermore, we conduct ablation studies to investigate the key factors influencing model performance.
4.1 Settings
Our evaluation encompassed three distinct types of models. Firstly, we examined OLMs, including several open-source contenders like VideoLLaMA 2 [cheng_videollama_2024], Unified-IO 2 [lu2023unifiedio2scalingautoregressive], Qwen2.5-Omni [xu_qwen25-omni_2025], Ola [liu_ola_2025] and our Daily-Omni Agent, as well as the proprietary Gemini models [noauthor_introducing_2024]. Secondly, we assessed VLMs, represented by the open-source Qwen2.5-VL [bai_qwen25-vl_2025] and the proprietary GPT-4o [openai_gpt-4o_2024]. We also evaluated the performance of some OLMs when provided with only visual inputs. Finally, we tested some LLMs such as Deepseek-V3 [deepseekai2025deepseekv3technicalreport], GPT-4o [openai_gpt-4o_2024] and Qwen2.5-14B[qwen2025qwen25technicalreport] without visual and audio inputs to check whether our benchmark contains too much information in questions and choices. All model evaluations were conducted strictly according to instructions from their respective official repositories and accompanying documentation (such as "cookbooks" or developer guides).
4.2 Main Results
Table 2 presents comprehensive evaluation results, shedding light on the real-world audio-visual understanding capabilities of contemporary MLLMs and our proposed Daily-Omni Agent. Firstly, earlier Omni-modal Language Models (OLMs) such as Unified-IO 2 and VideoLLaMA 2 demonstrate limited performance on our benchmarks. Notably, their results are, in several instances, even surpassed by text-only LLMs. Furthermore, the Unified-IO 2 series displays a perplexing degradation in performance with increasing model size, an observation consistent with findings from OmniBench [li_omnibench_2024]. A plausible explanation is an inadequate capacity within these language models to effectively process and synthesize cross-modal information.
Secondly, while overall performance still presents challenges, more recent open-source OLMs such as the Qwen2.5-Omni series and Ola demonstrate reasonable proficiency in inference and reasoning tasks. This suggests a foundational understanding of general visual and audio contexts. However, their efficacy significantly diminishes on temporally-sensitive tasks like audio-visual event alignment and context understanding, where results remain largely unsatisfactory. Despite incorporating cross-modal positional embeddings designed to provide temporal awareness, these models still struggle to address questions requiring precise temporal capabilities. Conversely, proprietary OLMs such as Gemini 2.0 Flash exhibit markedly superior cross-modal temporal capabilities, attaining the leading overall performance score of 67.84%. The observation that even top-tier OLMs have not yet breached the 70% accuracy threshold underscores the demanding nature of this benchmark, positing it as a valuable and pertinent objective for the advancement of contemporary OLMs.
Thirdly, upon ablating audio input and providing only visual data, all evaluated OLMs exhibit a substantial decline in performance. Notably, the magnitude of this performance loss tends to be greater for models that demonstrated superior initial performance when leveraging both modalities. The significant performance drop, especially in high-performing models, when deprived of audio, validates that the tasks within this benchmark genuinely require and benefit from the integration of both auditory and visual information, rather than being solvable by visual cues alone. This dependence on both modalities is further evidenced by the performance of leading Visual Language Models (VLMs). For instance, Qwen2.5-VL (7B) achieves only 40.68% accuracy, and even the highly capable GPT-4o (with its visual input) reaches just 56.47%. Given that these are state-of-the-art models within their respective size categories (sub-7B and larger proprietary models), their comparatively modest performance on our benchmark further underscores its value in demanding genuine audio-visual integration, a capability absent in previous multi-modal benchmarks.
These results reveal that: (1) Current OLMs demonstrate reasonable general audio and visual understanding capabilities when processing videos under 60 seconds in duration. However, they continue to struggle with tasks demanding sophisticated cross-modal temporal awareness or the integration of information across extended time intervals. (2) Our proposed question-answer generation pipeline proves effective in creating challenging evaluation instances that accurately probe the audio-visual understanding capabilities of MLLMs.
Table 2: Performance comparison of various MLLMs across different question domains and video durations. For Omni-modal Language Models (OLMs) with audio and visual inputs, Boldface shows the best performance and underlines show the second best. For OLMs with visual inputs only, subscripts on the ’Avg’ score demonstrate the performance difference compared to their audio-visual counterparts. For your reference, the accuracy rate of pure guessing is 25%. More details of model evaluation are provided in Appendix C.
Methods AV Event Alignment Comparative Context Understanding Event Sequence Inference Reasoning 30s Subset 60s Subset Avg
Omni-Modal Language Models (With Visual and Audio)
Qwen2.5-Omni (7B) $44.12$ $51.15$ $38.86$ $40.52$ $57.79$ $61.71$ $46.68$ $48.36$ $47.45$
Qwen2.5-Omni (3B) $38.66$ $48.09$ $33.68$ $33.99$ $54.55$ $44.00$ $42.35$ $38.36$ $40.52$
Ola (7B) $40.34$ $61.07$ $40.41$ $43.46$ $63.64$ $69.71$ $51.47$ $49.82$ $50.71$
Unified-IO-2 L (1B) $27.31$ $22.90$ $26.42$ $27.78$ $29.87$ $29.14$ $27.67$ $27.09$ $27.40$
Unified-IO-2 XL (3B) $30.25$ $30.53$ $25.39$ $29.08$ $33.12$ $21.71$ $28.13$ $28.55$ $28.32$
Unified-IO-2 XXL (8B) $25.63$ $31.30$ $26.42$ $25.82$ $35.06$ $29.71$ $26.74$ $30.00$ $28.24$
VideoLLaMA2 (7B) $35.71$ $35.88$ $35.75$ $31.70$ $40.91$ $34.29$ $38.02$ $31.82$ $35.17$
Gemini 2.0 Flash 62.18 73.28 63.73 63.72 76.62 75.43 67.23 68.55 67.84
Gemini 2.0 Flash Lite 55.04 $64.89$ $58.03$ 54.25 $74.03$ 72.00 $62.44$ 60.00 $61.32$
Daily-Omni (ours) $51.68$ 68.70 60.10 $53.92$ 78.57 $71.43$ 63.99 $59.27$ 61.82
Omni-Modal Language Models (Visual Only)
Qwen2.5-Omni (7B) $38.24$ $48.85$ $34.72$ $36.27$ $51.95$ $45.71$ $40.80$ $41.64$ $41.19_{\text{-}6.3}$
Qwen2.5-Omni (3B) $33.61$ $42.75$ $36.27$ $33.33$ $49.35$ $38.86$ $38.79$ $36.55$ $37.76_{\text{-}2.8}$
Gemini 2.0 Flash 39.08 64.12 56.48 56.21 67.53 $62.29$ $56.57$ $55.45$ 56.06-${}_{\text{-}11.8}$
Gemini 2.0 Flash Lite $43.70$ $58.02$ $53.89$ $45.10$ $64.29$ $60.57$ $53.01$ $51.64$ 52.38-${}_{\text{-}8.9}$
Visual Language Models (Visual Only)
GPT-4o $47.90$ $62.60$ $52.33$ $52.61$ 66.23 66.29 55.64 57.45 56.47
Qwen2.5-VL (7B) 36.97 46.56 $33.68$ 37.91 51.95 44.00 39.26 42.36 40.68
Qwen2.5-VL (3B) $35.71$ $43.51$ 34.72 $33.66$ $43.51$ $39.43$ $37.71$ $37.09$ $37.43$
Textual Language Models (Without Visual and Audio)
GPT-4o $33.19$ $43.51$ $28.50$ $30.39$ $44.81$ $46.86$ $36.48$ $36.18$ $36.34$
Deepseek-V3 (671B) $31.93$ $41.22$ $29.02$ $29.41$ $44.81$ $46.29$ $35.24$ $36.00$ $35.59$
Qwen2.5-Instruct (14B) $30.25$ $39.69$ $27.98$ $28.43$ $42.21$ $42.86$ $32.15$ $35.82$ $33.83$
4.3 Results of Daily-Omni Agent
Our Daily-Omni Agent, leveraging Qwen2.5-VL-7B, Qwen2-Audio (7B), and Qwen2.5-Instruct-14B, achieves 61.82% overall accuracy. This performance is state-of-the-art among open-source methods and surpasses smaller proprietary MLLMs. The success of our agent suggests that by simply leveraging VLM and ALM with time-period-level alignment coupled with event-level alignment for some key events, we are able to create a very strong OLM.
To further study the effect of event alignment, we conducted an ablation study evaluating the Daily-Omni Agent’s performance under three alignment scenarios: (1) No Alignment: Generating comprehensive visual and audio captions as a base, but omitting any form of event alignment. (2) Naive Alignment: Generating comprehensive visual and audio captions, followed by a step where the VLM is prompted to extract question-relevant events and their temporal boundaries (begin/end times). The ALM is then utilized to produce audio captions specifically for those extracted segments and attaining aligned audio-visual event pairs in the process. (3) Smart Alignment: Generating comprehensive visual and audio captions and aligned events as stated in Section 3.3. Note that for (2) and (3), we only consider generating an align event pair when the provided segment has a duration below a certain threshold to prevent confusion. The overall accuracy and the average number of aligned event pairs identified per question for each method are presented in Table 3. As expected, the Smart Alignment method achieves the highest average accuracy. This demonstrates that explicitly identifying and aligning relevant events significantly boosts the model’s overall performance compared to simply generating global captions. Conversely, the Naive Alignment method exhibited a slight decline in accuracy relative to the No Alignment baseline. Careful analysis of the generated events and their temporal boundaries suggests that this outcome is likely attributable to the limitations of the Qwen2.5-VL-7B model. It appears that this model is not sufficiently powerful to reliably identify and retrieve the target event through a single query. Moreover, the temporal grounding process itself frequently yields imprecise results, leading to erroneous alignment. This issue with temporal grounding affects even the Smart Alignment method, occasionally producing incorrect or confusing aligned event pairs. Consequently, we hypothesize that equipping the agent with a more powerful open-vocabulary video temporal grounding model would unlock further significant improvements in its performance, mitigating the impact of imprecise temporal grounding.
Figure 6: MLLMs’ accuracy over different question categories
Align Method Average Accuracy Aligned Event Pairs per Question
No Alignment $60.65$ 0
Naive Alignment $59.65$ 0.68
Smart Alignment 61.82 1.11
Table 3: Impact of different aligning methods on average accuracy and aligned event pairs per question.
4.4 Case Study
This part delves into selected case studies of model performance on Daily-Omni. Such an analysis of specific examples offers deeper insights into both the nature of the Daily-Omni benchmark and the current capabilities and limitations of the models. Figure 7 presents three illustrative questions from the Daily-Omni benchmark. It displays the responses of the Gemini 2.0 Flash and Qwen2.5-Omni models under two conditions: processing full audio-visual inputs versus visual-only inputs.
Figure 7: OLM responses to Daily-Omni questions: A comparison of audio-visual versus visual-only inputs. The figure presents examples of OLM performance when using full audio-visual modalities compared to visual-only input. For each case, it shows the audio-visual content, the question with choices (correct answer highlighted), and the models’ answers, indicating correctness for each input condition.
Case 1: Chronological Event Ordering This question requires determining the first event to occur in the video from a given list of audio-visual events. To correctly ascertain the existence and chronological position of choices B (Player discussing a personal milestone) and D (Female presenter announcing a groundbreaking sports contract), the model must leverage both audio and visual sensory inputs. Gemini 2.0 Flash successfully identifies choice B as the initial event, whereas Qwen2.5-Omni 7B fails to do so. Both model failed to answer correctly when only taking visual input. This disparity suggests that merely possessing multi-modal sensory capabilities is insufficient; models also require robust cross-modal reasoning to effectively integrate audio-visual information and achieve correct temporal ordering.
Case 2: Audio-Visual Inference This case assesses the ability to discern the functional differences between the hand gestures of two individuals engaged in a discussion. While visual input captures the gestures themselves, audio input is also critical for interpreting their communicative purpose as it reveals what is being said and who said it. Gemini 2.0 Flash correctly identified the functional difference (D) with audio-visual input but failed with visual-only input, selecting a more generic but incorrect interpretation (A). Qwen2.5-Omni 7B, failed to make the correct distinction in both audio-visual and visual-only conditions, selecting an unreasonable option (B). This question underscores that merely understanding the speech content is insufficient; accurately identifying the speaker for each utterance is also crucial. Successfully attributing speech requires robust audio-visual temporal integration—the ability to precisely synchronize the spoken audio with the visual depiction of the active speaker. Following this, strong reasoning ability is needed to infer the distinct roles and communicative intents (e.g., inquirer vs. narrator) based on the combined audio-visual information, and thereby understand the different functions of their gestures. The failure of Qwen2.5-Omni 7B, even with audio-visual input, suggests potential limitations in one or both of these sophisticated cross-modal capabilities.
Case 3: Sound Understanding This question directly tests the model’s ability to analyze and compare audio characteristics across two different segments of a video-an skateboard demonstration and an outdoor skatepark scene. The task involves identifying distinct audio elements such as human speech, specific sound effects (e.g., skateboard impact), and background music. This is fundamentally an audio-centric task where visual context primarily helps situate the sounds. Strikingly, both Gemini 2.0 Flash and Qwen2.5-Omni failed to correctly compare the audio content, selecting the same incorrect option (D) irrespective of whether they received audio-visual or visual-only input. Their failure in the audio-visual condition suggests a significant challenge in fine-grained auditory scene analysis, such as distinguishing non-verbal sound event and music, or accurately mapping these perceived audio features to the descriptive choices provided.
5 Conclusion
This paper introduced Daily-Omni, a novel Audio-Visual Question Answering benchmark designed to evaluate MLLMs on temporally-aligned multimodal reasoning in daily life scenarios. We also proposed an efficient Daily-Omni QA Generation Pipeline and the training-free Daily-Omni Agent, a strong open-source baseline. Our evaluation revealed that while recent MLLMs show general audio-visual understanding, they significantly struggle with precise cross-modal temporal awareness The results underscore our pipeline’s value and effectiveness in creating such challenging questions. The Daily-Omni Agent, with its targeted temporal alignment, achieved competitive results, and our ablation study confirmed that a Smart Alignment strategy improves performance, unlike naive approaches. In summary, Daily-Omni highlights the current MLLM frontiers in complex audio-visual temporal reasoning. Future efforts should prioritize developing more accurate and robust multimodal temporal grounding techniques for truly sophisticated audio-visual understanding.
Appendix A Societal Impact
The Daily-Omni benchmark and the associated research aim to advance the capabilities of Multimodal Large Language Models (MLLMs) in understanding and reasoning about real-world audio-visual scenarios.
Positive Societal Impacts:
  • Improved Human-AI Interaction: MLLMs with better audio-visual understanding can lead to more natural and effective human-AI interaction in various applications, such as assistive technologies for individuals with disabilities (e.g., describing surroundings for visually impaired individuals), smarter virtual assistants, and more intuitive robotics.
  • Enhanced Content Accessibility and Analysis: Advanced MLLMs can automatically generate rich descriptions, summaries, and answer questions about video content, making vast amounts of audio-visual data more accessible and searchable. This can benefit education, media archiving, and content moderation.
  • Development of Artificial Agent Systems capable of perceiving and understanding the outcome of their decision or operation via audio and visual modalities will be smarter.
Negative Societal Impacts and Mitigation:
  • Surveillance and Privacy Concerns: Enhanced audio-visual understanding capabilities could be deployed in surveillance systems, raising significant privacy concerns if not implemented with strong ethical guidelines and legal oversight. Mitigation: Development and deployment of such technologies must be accompanied by robust privacy-preserving techniques (e.g., federated learning, on-device processing, data anonymization) and strong regulatory frameworks that protect individual privacy. The videos in Daily-Omni are sourced from publicly available platforms, but the broader application of the technology requires careful consideration.
  • Bias Amplification: MLLMs are trained on large datasets, which may contain societal biases. These biases can be learned and amplified by the models, leading to unfair or discriminatory outcomes in their applications. Mitigation: Efforts in dataset curation for Daily-Omni aimed for diversity in video categories (Section 3.1, Figure 2). Ongoing research into bias detection, fairness metrics, and debiasing techniques for MLLMs is essential. Continuous auditing of model performance across different demographic groups is also necessary.
Our work primarily focuses on creating a benchmark to evaluate and drive progress in MLLM capabilities. We believe that by transparently assessing these models, we can better understand their strengths and weaknesses, which is a prerequisite for responsible development and deployment.
Appendix B Licenses for Existing Assets
This research utilizes several existing assets, including datasets and pre-trained models. We have made efforts to properly credit the original creators through citations in the main paper and adhere to their terms of use.
Datasets Used for Video Sourcing:
  • AudioSet [Audioset]: The AudioSet dataset consists of an expanding ontology of 632 audio event classes and a collection of 2,084,320 human-labeled 10-second sound clips drawn from YouTube videos. The video segments used in Daily-Omni were derived from the original full-length YouTube videos corresponding to AudioSet clips. AudioSet itself is released by Google under a Creative Commons license (typically CC BY 4.0 for the annotations, while the YouTube videos are subject to their own licenses set by uploaders). We accessed publicly available YouTube videos.
  • Video-MME [fu2024videommefirstevercomprehensiveevaluation]: Video-MME is a benchmark for evaluating multimodal models on video understanding. The videos are sourced from YouTube. The benchmark is only used for academic research. Commercial use in any form is prohibited.
  • FineVideo [Farré2024FineVideo]: FineVideo is a dataset for fine-grained video content analysis. The videos provided are derived from YouTube-Commons. All the videos are shared under a CC-By license and, in accordance with that license, every YouTube channel is fully credited.
For all video data sourced from YouTube, we rely on the videos being publicly accessible and adhere to YouTube’s Terms of Service. The specific license for each YouTube video is determined by its uploader. Our use is for non-commercial research purposes to create a benchmark.
Models Used (Annotation, QA Generation, Daily-Omni Agent, Evaluation):
  • Whisper [radford2022robustspeechrecognitionlargescale]: Developed by OpenAI. The Whisper models are typically released under a permissive license (e.g., MIT License), allowing for research and commercial use. We used Whisper-Large-V2 and Whisper-V3-Large.
  • Gemini Models [noauthor_introducing_2024]: Developed by Google. Access to Gemini models (e.g., Gemini 2.0 Flash, Gemini 2.0 Flash Lite) is typically through Google’s APIs (e.g., Google AI Studio or Vertex AI), subject to Google’s terms of service and API usage policies.
  • Deepseek Models (Deepseek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability], Deepseek-V3 [deepseekai2025deepseekv3technicalreport]): Developed by DeepSeek AI. These models are often available via API or with specific research/community licenses. Their terms of use were respected.
  • Qwen Models (Qwen2-Audio [chu2024qwen2audiotechnicalreport], Qwen2.5-VL [bai_qwen25-vl_2025], Qwen2.5-Omni [xu_qwen25-omni_2025], Qwen2.5-14B-Instruct [qwen2025qwen25technicalreport]): Developed by Alibaba Cloud. Qwen models are often released under licenses like the Tongyi Qianwen LICENSE AGREEMENT, which may have specific restrictions for commercial use but generally permit research.
  • GPT-4o [openai_gpt-4o_2024]: Developed by OpenAI. Accessed via OpenAI’s API, subject to their terms of service and usage policies.
  • Other MLLMs for Evaluation (VideoLLaMA 2 [cheng_videollama_2024], Unified-IO 2 [lu2023unifiedio2scalingautoregressive], Ola [liu_ola_2025]): These models are from various research institutions and companies. Their use in our evaluation adheres to the licenses provided with their respective releases (e.g., Apache 2.0, MIT, or custom research licenses). Original papers are cited.
We have strived to use all assets in accordance with their intended purpose for research and benchmarking, and have credited original authors. The Daily-Omni benchmark itself, including the QA pairs, will be released under CC BY-NC-SA 4.0 to encourage its use and further development in the research community, while respecting the licenses of the underlying video content.
Appendix C Evaluation Details
The models listed in Table 2 were evaluated as follows: Qwen2.5-Omni (7B and 3B), Ola, Unified-IO-2 (L, XL, and XXL), VideoLLaMA2, and Qwen2.5-VL (7B and 3B) were deployed and tested on a local server. In contrast, Gemini 2.0 Flash, Gemini 2.0 Flash Lite, GPT-4o, and Qwen2.5-Instruct (14B) were evaluated via their respective APIs.
For the Daily-Omni Agent, Qwen2.5-VL-7B-Instruct and Qwen2.5-14B-Instruct were accessed via the API provided by Alibaba Cloud’s Bailian platform (accessible at https://bailian.console.aliyun.com/) to ensure operational efficiency. Similarly, Whisper-Large-V2 was utilized via the API provided by OpenAI. The Qwen2-Audio component, on the other hand, was deployed and tested locally. It is important to note that all component models of Daily-Omni are capable of local deployment. According to official documentation from Aliyun (https://help.aliyun.com/zh/model-studio/model-user-guide/) and OpenAI (https://openai.com/index/introducing-chatgpt-and-whisper-apis/), the Qwen models offered through Bailian and the Whisper-Large-V2 model accessed via the OpenAI API are identical to their respective open-source or standard versions. Therefore, no performance differences are anticipated.
Our code implements direct passing of local video path to the Qwen2.5-VL API. However, this functionality might require you to contact Aliyun customer service to enable. If direct path input is not activated, you can alternatively pass a list of video frames, though this may result in suboptimal performance.
[…]

### v2

@@
 However, the ability of these models to process cross-modal information synchronously remains largely unexplored.
-In this paper, we introduce: 1) Daily-Omni, an Audio-Visual Questioning and Answering benchmark comprising 684 videos of daily life scenarios from diverse sources, rich in both audio and visual information, and featuring 1197 multiple-choice QA pairs across 6 major tasks; 2) Daily-Omni QA Generation Pipeline, which includes automatic annotation, QA generation and QA optimization, significantly improves efficiency for human evaluation and scalability of the benchmark; 3) Daily-Omni-Agent, a training-free agent utilizing open-source Visual Language Model (VLM), Audio Language Model (ALM) and Automatic Speech Recognition (ASR) model to establish a baseline for this benchmark.
-The results show that current MLLMs still struggle significantly with tasks requiring audio-visual integration, but combining VLMs and ALMs with simple temporal alignment techniques can achieve substantially better performance.
-Codes and benchmark are available at https://github.com/Lliar-liar/Daily-Omni.
+We introduce Daily-Omni, a multiple-choice Audio-Visual QA benchmark featuring 684 real-world videos and 1,197 questions spanning 6 task families that explicitly require cross-modal temporal reasoning.
+To support scalable benchmark construction, we develop a semi-automatic pipeline for annotation, cross-modal consistency refinement, temporal alignment elicitation, and text-only leakage filtering, followed by human verification.
+We further provide a diagnostic evaluation suite and extensively evaluate 24 foundation models under 37 model–modality settings (Audio+Video / Audio-only / Video-only / Text-only).
+Finally, we include a training-free modular diagnostic baseline that composes off-the-shelf unimodal models to serve as a diagnostic baseline and to illustrate how explicit temporal alignment signals affect performance.
+Results indicate that many end-to-end MLLMs still struggle on alignment-critical questions, suggesting that robust cross-modal temporal alignment remains an important open challenge.
 1 Introduction
-Non-textual modalities, such as vision and hearing, often convey richer and more nuanced information than text alone in daily life scenarios, playing a crucial role in our understanding of and interaction with the physical world.
+In our daily lives, non-textual modalities like vision and audio convey much richer information than text alone.
+They are essential for us to understand and interact with the physical world.
 Therefore, advancements in Multimodal Large Language Models (MLLMs) capable of comprehensively understanding the multi-modal information are a crucial foundation for achieving artificial intelligence capable of interacting with the physical world and observing the outcome of its operations.
-Recent MLLMs [xu_qwen25-omni_2025, sun_video-salmonn_2024, noauthor_introducing_2024, zhang_internlm-xcomposer25-omnilive_2024, fu_vita-15_2025, liu_ola_2025, cheng_videollama_2024, li_baichuan-omni-15_2025, xie_mini-omni2_2024, guo_m2-omni_2025, team_reka_2024, microsoft2025phi4minitechnicalreportcompact, lu2023unifiedio2scalingautoregressive, liu_nexus-o_2025, gemmateam2025gemma3technicalreport] have demonstrated groundbreaking capabilities spanning audio (ASR, sound classification, captioning) and visual domains (OCR, VQA, video grounding), with significant accuracy improvements over previous benchmarks.
+Recent MLLMs [Qwen3-VL, Qwen3-Omni, xu_qwen25-omni_2025, tang_enhancing_2024, noauthor_introducing_2024, zhang_internlm-xcomposer25-omnilive_2024, fu_vita-15_2025, liu_ola_2025, cheng_videollama_2024, li_baichuan-omni-15_2025, xie_mini-omni2_2024, guo_m2-omni_2025, team_reka_2024, microsoft2025phi4minitechnicalreportcompact, lu2023unifiedio2scalingautoregressive, liu_nexus-o_2025, gemmateam2025gemma3technicalreport, gemini25] have demonstrated groundbreaking capabilities spanning audio (ASR, sound classification, captioning) and visual domains (OCR, VQA, video grounding), with significant accuracy improvements over previous benchmarks.
 However, existing methods still face several limitations.
@@
 Secondly, the current landscape lacks high-quality multimodal datasets that integrate temporally aligned auditory and visual information.
-Existing audio-visual datasets and benchmarks [yun_pano-avqa_2021, li_learning_2022, avqa, li_omnibench_2024, hong_worldsense_2025, gong_av-odyssey_2024, sungbin2025avhbenchcrossmodalhallucinationbenchmark, geng2025longvalevisionaudiolanguageeventbenchmarktimeaware] reveal three persistent limitations.
-First, several focus on specialized scenarios [yun_pano-avqa_2021, li_learning_2022] such as musical performances or panoramic environments, thereby introducing domain-specific biases.
+Existing audio-visual datasets and benchmarks [yun_pano-avqa_2021, li2022learninganswerquestionsdynamic, avqa, li_omnibench_2024, hong_worldsense_2025, gong_av-odyssey_2024, sungbin2025avhbenchcrossmodalhallucinationbenchmark, geng2025longvalevisionaudiolanguageeventbenchmarktimeaware, omnivideobench] suffer from three persistent limitations.
+First, several focus on specialized scenarios [yun_pano-avqa_2021, li2022learninganswerquestionsdynamic] such as musical performances or panoramic environments, thereby introducing domain-specific biases.
+Figure 1: Illustrative examples from the Daily-Omni benchmark.
+These cases highlight the dataset’s focus on complex cross-modal reasoning.
+The top example requires temporally aligning specific speakers’ voices with their gestures, while the bottom example requires comparing acoustic environments across different video segments.
 Second, many employ static image-audio pairs [li_omnibench_2024, gong_av-odyssey_2024] that disregard crucial temporal dynamics inherent in real-world video contexts.
@@
 This dual limitation constrains both the benchmark’s adaptability to emerging domains and its practical utility in driving model improvements.
-Figure 1: Examples of Daily-Omni QAs. The audio and visual information required for answering the questions are provided in the figure. The correct answer for the given questions are highlighted.
-In this paper, we introduce Daily-Omni, an Audio-Visual Questioning and Answering benchmark with 684 videos featuring daily life scenarios from various sources [Audioset, fu2024videommefirstevercomprehensiveevaluation, Farré2024FineVideo] with rich audio and visual information and 1197 multiple choice QA-pairs across 6 major tasks ranging from audio visual event aligning to complicated cross-modal reasoning.
-The videos are sampled from different datasets and segmented into 30-second or 60-second intervals to systematically evaluate model performance across different temporal contexts.
-We further introduce Daily-Omni QA Generation Pipeline which encompasses five automated modules: video annotation, annotation revision, audio-visual temporal alignment, QA generation, and QA optimization.
-This framework demonstrates remarkable scalability: a single annotator can complete quality filtering process for these 1197 high-quality QA pairs within 30 hours, achieving an approximate 30% acceptance rate from initially generated candidates.
-In addition, we propose Daily-Omni Agent, an audio-visual agent utilizing open-source Visual Language Model (VLM), Audio Language Model (ALM) and Automatic Speech Recognition (ASR) model without further finetuning to establish a baseline for this benchmark.
-We evaluated recent MLLMs and our agent on the Daily-Omni benchmark, where the Daily-Omni Agent achieved state-of-the-art performance among open-source methods.
-The experimental results show that current MLLMs still face significant challenges in tasks requiring deep audio-visual temporal integration.
-They also reveal that by combining existing visual and audio language models with simple temporal alignment techniques, as demonstrated by our Daily-Omni Agent, substantially improved performance can be achieved, underscoring a promising direction for enhancing multimodal reasoning.
+In this paper, we introduce Daily-Omni, an Audio-Visual Questioning and Answering benchmark designed to evaluate temporally-aligned multimodal reasoning.
+The videos are sampled from diverse datasets [Audioset, fu2024videommefirstevercomprehensiveevaluation, Farre2024FineVideo] and segmented into 30-second or 60-second intervals to systematically assess model performance across different temporal contexts.
+Our core contributions are three-fold:
+  • The Daily-Omni Benchmark: We construct a high-quality dataset comprising 684 real-life videos rich in both audio and visual dynamics, featuring 1197 multiple-choice QA pairs across 6 major tasks ranging from fundamental audio-visual event alignment to complicated cross-modal reasoning.
+  • Scalable QA Generation Pipeline: We propose an automated framework encompassing video annotation, revision, audio-visual temporal alignment, QA generation, and optimization.
+This highly efficient pipeline significantly reduces human labor, requiring only 30 hours for a single annotator to complete the quality filtering process (achieving an approximate 30% acceptance rate), thereby ensuring the benchmark’s scalability.
+  • Extensive Evaluation & Diagnostic Protocol.
+We benchmark 24 representative foundation models under 37 model–modality settings (AV / V-only / A-only / text-only) and provide diagnostic analyses including modality ablations and alignment-sensitivity.
+We additionally include a training-free modular reference baseline (Daily-Omni Agent) composed of off-the-shelf unimodal models, intended as a sanity-check reference.
+Our comprehensive results reveal an important phenomenon: many current end-to-end MLLMs struggle with tasks requiring deep audio-visual temporal integration, to the extent that our simple, explicit-alignment baseline outperforms several recent open-source omni-modal models.
+This extensive evaluation demonstrates that robust temporal alignment mechanisms are still insufficient or missing in many unified architectures, underscoring both the necessity of our benchmark and a clear direction for future multimodal reasoning research.
 2 Related Works
 2.1 Multimodal Large Language Models
-Recent advancements in Multimodal Large Language Models (MLLMs) primarily fall into three categories: Audio Language Models (ALMs) incorporating audio capabilities [tang_salmonn_2024, gong2024listenthinkunderstand, chu_qwen-audio_2023, chu2024qwen2audiotechnicalreport, ghosh2025audioflamingo2audiolanguage, ghosh2024gamalargeaudiolanguagemodel], Visual Language Models (VLMs) adding visual understanding [liu2025nvilaefficientfrontiervisual, bai_qwen25-vl_2025, lu2024deepseekvlrealworldvisionlanguageunderstanding, wang2024emu3nexttokenpredictionneed], and Omni-modal Language Models (OLMs) combining both audio and visual modalities [noauthor_introducing_2024, li_baichuan-omni-15_2025, zhang_internlm-xcomposer25-omnilive_2024, team_reka_2024, cheng_videollama_2024, liu_ola_2025, microsoft2025phi4minitechnicalreportcompact, fu_vita-15_2025, guo_m2-omni_2025, xu_qwen25-omni_2025, liu_nexus-o_2025, gemmateam2025gemma3technicalreport].
+Recent advancements in Multimodal Large Language Models (MLLMs) primarily fall into three categories: Audio Language Models (ALMs) incorporating audio capabilities [tang_salmonn_2024, gong2024listenthinkunderstand, chu_qwen-audio_2023, chu2024qwen2audiotechnicalreport, ghosh2025audioflamingo2audiolanguage, ghosh2024gamalargeaudiolanguagemodel, goel2025audioflamingo3advancing], Visual Language Models (VLMs) adding visual understanding [liu2025nvilaefficientfrontiervisual, bai_qwen25-vl_2025, lu2024deepseekvlrealworldvisionlanguageunderstanding, wang2024emu3nexttokenpredictionneed], and Omni-modal Language Models (OLMs) combining both audio and visual modalities [noauthor_introducing_2024, li_baichuan-omni-15_2025, zhang_internlm-xcomposer25-omnilive_2024, team_reka_2024, cheng_videollama_2024, liu_ola_2025, microsoft2025phi4minitechnicalreportcompact, fu_vita-15_2025, guo_m2-omni_2025, Qwen3-Omni, xu_qwen25-omni_2025, liu_nexus-o_2025, gemmateam2025gemma3technicalreport].
 These MLLMs often employ a modular architecture, utilizing separate encoders for audio and visual inputs.
-In the audio domain, encoders such as [radford2022robustspeechrecognitionlargescale, chen_beats_2022, li2024mertacousticmusicunderstanding] are commonly used to extract sound representations.
-Some models, such as [tang_salmonn_2024, liu_ola_2025, zhang_internlm-xcomposer25-omnilive_2024], even integrate multiple audio encoders specialized for different sound types (e.g., speech, music).
-Similarly, visual encoders [liu2021swintransformerhierarchicalvision, li2023blip2bootstrappinglanguageimagepretraining, dai2023instructblipgeneralpurposevisionlanguagemodels, liu2023visualinstructiontuning, dehghani2023patchnpacknavit] are used to process images and videos.
-However, this modular approach, while cost-efficient and effective for individual modalities, struggles to capture the crucial temporal correlations inherent in synchronized audio-visual streams.
+In the audio domain, several works [chen2022beatsaudiopretrainingacoustic, li2024mertacousticmusicunderstanding, radford2022robustspeechrecognitionlargescale] proposed audio encoders to extract sound representations.
+Some models [tang_salmonn_2024, liu_ola_2025, zhang_internlm-xcomposer25-omnilive_2024] even integrate multiple audio encoders specialized for different sound types (e.g., speech, music).
+Similarly, visual encoders [li2023blip2bootstrappinglanguageimagepretraining, dai2023instructblipgeneralpurposevisionlanguagemodels, liu2023visualinstructiontuning, dehghani2023patchnpacknavit] are used to process images and videos.
+However, this modular approach struggles to capture the crucial temporal correlations inherent in synchronized audio-visual streams.
 This limitation arises because audio and visual inputs are encoded independently.
-Although multi-modal positional embedding techniques like M-RoPE [bai2023qwenvlversatilevisionlanguagemodel] and TMRoPE [xu_qwen25-omni_2025] enhance cross-modal temporal understanding to some extent, effective methods for deeply integrating and temporally aligning multimodal data remain relatively scarce.
-Additionally, in several MLLMs [microsoft2025phi4minitechnicalreportcompact, fu_vita-15_2025, zhang_internlm-xcomposer25-omnilive_2024] , the audio encoder primarily serves to process user instructions–akin to how the text modality is used–rather than sensing the ambient environment, which is a common function of the visual modality.
+Although multi-modal positional embedding techniques like TMRoPE [xu_qwen25-omni_2025] enhance cross-modal temporal understanding to some extent, effective methods for temporally aligning multimodal data remain relatively scarce.
+Additionally, in several MLLMs [microsoft2025phi4minitechnicalreportcompact, fu_vita-15_2025, zhang_internlm-xcomposer25-omnilive_2024] , the audio encoder primarily serves to process user instructions—akin to how the text modality is used—rather than perceiving the environment.
+Table 1: Comparison of audio-visual benchmarks.
+We detail publication, modality (A: audio, V: video, I: image), size, and question type (MC: multiple choice, DF: defined word, $B$: bounding boxes).
+’Automatic Scalability’ refers to automated expansion methods.
+’Open-Domain’ and ’General Sound’ denote genre and sound diversity.
+Benchmarks Pub Modality #QA Question Type Automatic Scalability Open-Domain General Sound
+AVQA ACM MM’22 V+A 57,335 MC
+Music-AVQA CVPR’22 V+A 45,867 DF
+Pano-AVQA ICCV’21 V+A 51,700 DF & BB
+OmniBench ARXIV’24 I+A 1,142 MC
+AV-Odyssey ARXIV’24 I+A 4,555 MC
+WorldSense ICLR’26 V+A 3,172 MC
+Daily-Omni V+A 1197 MC
 2.2 Audio-Visual Understanding Datasets and Benchmarks
 The development of uni-modal datasets and associated tasks for audio and vision has driven advancements in Audio Language Models (ALMs) and Visual Language Models (VLMs).
-Visual datasets and benchmarks [li2024llavanextinterleavetacklingmultiimagevideo, fu2024videommefirstevercomprehensiveevaluation, wu_star_2024, mangalam_egoschema_2023, liu2023visualinstructiontuning, wu_longvideobench_nodate, zhang2024videoinstructiontuningsynthetic, yue2024mmmumassivemultidisciplinemultimodal, yue2024mmmuprorobustmultidisciplinemultimodal, li2023seedbenchbenchmarkingmultimodalllms, fu2024ocrbenchv2improvedbenchmark, gao2017talltemporalactivitylocalization, hu2025videommmuevaluatingknowledgeacquisition] primarily focus on tasks for static images (such as OCR, grounding, segmentation, classification, and question-answering) and dynamic videos (such as captioning, temporal grounding, and understanding), while audio datasets and benchmarks [ghosh2025audioflamingo2audiolanguage, Audioset, chen_vggsound_2020, Librispeech, yang2024airbenchbenchmarkinglargeaudiolanguage] address speech-related tasks (including ASR, emotion recognition, and entity recognition) and non-speech tasks (such as sound classification and audio grounding).
+Visual datasets and benchmarks [li2024llavanextinterleavetacklingmultiimagevideo, fu2024videommefirstevercomprehensiveevaluation, wu_star_2024, mangalam_egoschema_2023, liu2023visualinstructiontuning, wu2024longvideobench, zhang2024videoinstructiontuningsynthetic, yue2024mmmumassivemultidisciplinemultimodal, yue2024mmmuprorobustmultidisciplinemultimodal, li2023seedbenchbenchmarkingmultimodalllms, fu2024ocrbenchv2improvedbenchmark, gao2017talltemporalactivitylocalization, hu2025videommmuevaluatingknowledgeacquisition, omnivideobench] primarily focus on tasks for static images (OCR, grounding, segmentation, classification, and question-answering) and dynamic videos (captioning, temporal grounding, and understanding), while audio datasets and benchmarks [ghosh2025audioflamingo2audiolanguage, Audioset, chen_vggsound_2020, Librispeech, yang2024airbenchbenchmarkinglargeaudiolanguage] address speech-related tasks (ASR, emotion recognition, and entity recognition) and non-speech tasks (such as sound classification and audio grounding).
 However, while attempts at creating audio-visual datasets date back to at least 2021 [yun_pano-avqa_2021], these early efforts often had significant limitations.
@@
 Furthermore, AV-Odyssey [gong_av-odyssey_2024] heavily emphasized specific audio tasks, such as recognizing timbre and loudness, rather than broader audio-visual understanding.
-While WorldSense [hong_worldsense_2025] provides a valuable model for real-world audio-visual question-answering, how to efficiently scale AVQA datasets and enhance OLM abilities is an area requiring further exploration.
+While WorldSense [hong_worldsense_2025], a concurrent work, also provides a valuable benchmark for real-world audio-visual question-answering, the efficient scalability of AVQA datasets and the enhancement of OLM abilities still require further exploration.
 3 Daily-Omni
-This section details the Daily-Omni framework, which comprises three core components: multi-modal data curation, hierarchical video annotation, and QA synthesis and evaluation.
-Furthermore, we delineate the evaluation paradigm implemented through the Daily-Omni Agent, which serves as our benchmarking baseline for systematic assessment.
-Daily-Omni aims to establish a benchmarking framework designed to systematically evaluate MLLMs’ ability to perceive and reason in real-life audio-visual scenarios.
-As illustrated in Figure LABEL:Fig:qa_count, the Daily-Omni benchmark comprises videos from all 11 YouTube categories.
-The QAs are meticulously crafted to assess a spectrum of multi-modal capabilities, spanning from basic cross-modal perception to intricate reasoning.
-All questions are designed to require the integration of audio, visual, and textual information for a correct answer.
-In total, the Daily-Omni benchmark comprises 684 videos and 1197 QA pairs.
-Of these QAs, 550 correspond to 60-second videos and 647 correspond to 30-second videos.
-We compared the features of Daily-Omni with various audio-visual benchmarks in Table 1.
+This section describes the construction of Daily-Omni, including data curation, video annotation, and QA synthesis and verification.
+We also describe an evaluation protocol and include a training-free modular reference baseline to help interpret model behaviors.
+Daily-Omni contains 684 real-world videos from all 11 YouTube categories and 1,197 multiple-choice questions spanning 6 task families.
+The dataset includes 550 questions on 60-second clips and 647 questions on 30-second clips.
+Table 1 compares Daily-Omni with prior audio-visual benchmarks.
+(a) Word cloud of Daily-Omni QA pairs
+(b) Statistics and distributions of the benchmark
+Figure 2: Overview of the Daily-Omni benchmark.
+(a) The prominent vocabulary in the word cloud highlights our focus on rich human activities, diverse acoustic events, and temporal reasoning.
+(b) The bar chart details the long-tail distribution of the 14 source video categories (top) and the composition of the 6 core question types (bottom).
 3.1 Data Curation
-Daily-life scenarios offer abundant information across both audio and visual modalities.
-Within the visual modality, our data curation focuses on selecting videos characterized by significant temporal dynamics.
-Conversely, static scenes, exemplified by a vlogger addressing the camera with limited motion, offer minimal temporal information for questioning and are usually excluded.
-To prevent data bias, videos are selected from a wide range of categories and topics.
-Videos with harmful or discriminative content are excluded.
-Regarding the audio modality, recognizing that prior datasets have concentrated heavily on speech, our benchmark is designed to encompass a broader range of everyday sounds, such as music, speech, and other sound events.
-These acoustic signals should be present within the videos, appearing either concurrently or consecutively.
-Furthermore, to maintain focus on audio-visual processing and avoid challenges related to multilingual text comprehension, videos with languages other than English are also excluded.
-Our video data originates from AudioSet [Audioset], Video-MME [fu2024videommefirstevercomprehensiveevaluation], and FineVideo [Farré2024FineVideo].
-While AudioSet provides 10-second clips primarily for audio classification, we retrieved the original, full-length source videos.
-These were then processed into longer 30s or 60s segments which contains the original 10s segments to ensure they contain certain types of sound events.
-To ensure inclusion of spoken content, we employed Whisper-V3-Large [radford2022robustspeechrecognitionlargescale] to detect speech within these segments, adding those segments identified as containing speech to our set of video candidates.
-For Video-MME and FineVideo, our selection criteria prioritized videos containing substantial audio information alongside rich temporal visual dynamics.
-The inclusion of these datasets also increases diversity, as they feature more recent videos compared to Audio Set (primarily pre-2017 uploads), thus broadening the representation of genres and styles.
+We curate videos with rich temporal dynamics in vision and diverse everyday sounds beyond speech (e.g., music and environmental events).
+We exclude largely static talking-head videos and non-English videos to reduce confounds from multilingual content.
+Figure 2 shows the word cloud of questions and choices in Daily-Omni and the distribution of video categories and QA types.
+Our videos are sourced from AudioSet [Audioset], Video-MME [fu2024videommefirstevercomprehensiveevaluation], and FineVideo [Farre2024FineVideo].
+For AudioSet, we retrieve the original YouTube videos and extract 30s/60s segments that contain the target sound events.
+We use Whisper-V3-Large [radford2022robustspeechrecognitionlargescale] to ensure spoken content is present.
+For Video-MME and FineVideo, we prioritize clips with substantial audio information and salient temporal visual changes to increase scenario diversity.
 3.2 Data Annotation & QA Construction
@@
 To obtain detailed annotations while ensuring cost-effectiveness, we specifically used Gemini 2.0 Flash [noauthor_introducing_2024] for the annotation task and Deepseek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability] for QA construction and optimization.
-Figure 3 provides a outline of this process.
+Figure 3 provides an outline of this process.
 Figure 3: The outline of Daily-Omni QA construction pipeline. The arrows indicates the sequence of the processes.
@@
 Visual & Audio Event Alignment At this stage, we have sequences of visual and audio events annotated within consecutive 10s or 20s segments.
-However, these segment-level annotations do not explicitly specify the temporal alignment between individual visual and audio events – i.e., which specific events occurred simultaneously.
-To establish this precise cross-modal event concurrency, we proposed a novel event aligning technique.
+However, these segment-level annotations do not explicitly specify the temporal alignment between individual visual and audio events, i.e., which specific events occurred simultaneously.
+To establish this precise cross-modal event concurrency, we proposed a event aligning technique.
 By prompting Gemini 2.0 Flash with the complete audio-visual clip, we instruct it to identify the visual event(s) occurring concurrently with each identified audio event.
@@
 The details of annotation generation, revision and event alignment is shown in Figure 4.
+We further assess the reliability of this alignment step via manual verification on a randomly sampled subset of 100 videos: over 90% of aligned audio–visual event pairs are judged correct.
 Figure 4: Details of Daily-Omni annotation generation, revision and event alignment. For cost-efficiency, we align all events with one query.
-QA Construction Using the consistent visual annotations, revised audio annotations, and aligned event pairs derived from each video, we prompted Deepseek-R1 to generate multi-choice questions covering the following types: (1) AV Event Alignment: Questions to determine which audio and visual events occurred simultaneously with each other; (2) Event Sequence: Questions to determine the temporal sequence of visual and audio events in the video; (3) Reasoning: Questions to explain the cause or reason behind the occurrence of a visual or audio event in the video; (4) Inference: Questions to speculate on information not explicitly presented in the video; (5) Comparative: Questions to compare the similarity or difference between the audio and visual information of two or more events in the video; (6) Context Understanding: Questions to determine the contextual information surrounding a specific event in the video.
+QA Construction Using the consistent visual annotations, revised audio annotations, and aligned event pairs derived from each video, we prompted Deepseek-R1 to generate multi-choice questions covering the following types: (1) AV Event Alignment: To determine which audio and visual events occurred simultaneously with each other; (2) Event Sequence: To determine the temporal sequence of visual and audio events in the video; (3) Reasoning: To explain the cause or reason behind the occurrence of a visual or audio event in the video; (4) Inference: To speculate on information not explicitly presented in the video; (5) Comparative: To compare the similarity or difference between the audio and visual information of two or more events in the video; (6) Context Understanding: To determine the contextual information surrounding a specific event in the video.
 Since Daily-Omni aims to evaluate MLLM perception and reasoning within real-life audio-visual scenarios, we deliberately exclude certain question types.
 Specifically, counting and measuring questions are omitted, as they often rely on a single modality rather than integrated multi-modal understanding.
-Similarly, purely knowledge-based queries, such as celebrity identification, fall outside our intended scope.
-Table 1: Comparison of various audio-visual benchmarks.
-The table details key characteristics including modality (A: audio, V: video, I: image), dataset size (#Video, #QA Pairs), and Question Type (MC: multiple choice, DF: defined word, BB: bounding boxes).
-’Efficient Scalability’ indicates whether the associated work proposed an automated method for benchmark expansion.
-’Open-Domain’ denotes benchmarks featuring various categories visual content, while ’General Sound’ signifies benchmarks whose audio content includes speech, music, and generic real-life sounds.
-Benchmarks Modality #Video #QA Pairs Question Type Efficient Scalability Open-Domain General Sound
-AVQA [avqa] V+A 57,015 57,335 MC
-Music-AVQA [li2022learninganswerquestionsdynamic] V+A 9,288 45,867 DF
-Pano-AVQA [yun_pano-avqa_2021] V+A 5,400 51,700 DF & BB
-OmniBench [li_omnibench_2024] I+A 1,142 MC
-AV-Odyssey [gong_av-odyssey_2024] I+A 4,555 MC
-WorldSense [hong_worldsense_2025] V+A 1,662 3,172 MC
-Daily-Omni (ours) V+A 684 1197 MC
+Similarly, purely knowledge-based queries such as celebrity identification are not included.
 QA Optimization and Quality Control By avoiding strict question templates, the generated QAs exhibit greater creativity, incorporate complicated logic, and sometimes feature obscure descriptions, making them more closely resemble questions asked in real-life scenarios.
@@
 A single annotator used less than 30 hours to review the candidates and establish the final set of 1197 QAs, corresponding to an acceptance rate of approximately 30% during this manual review stage.
-Figure 5: The outline of Daily-Omni Agent.
-3.3 Daily-Omni Agent
-To establish a baseline for MLLMs and demonstrate the importance of temporal awareness in audio-visual question answering, we constructed the "Daily-Omni Agent".
-This agent, designed to understand audio-visual context and answer related questions, integrates several models: Qwen2-Audio (7B parameters) [chu2024qwen2audiotechnicalreport], Qwen2.5-VL-7B [bai_qwen25-vl_2025], Whisper-Large-V2 [radford2022robustspeechrecognitionlargescale], and the text-based Qwen2.5-14B-Instruct [qwen2025qwen25technicalreport].
+Figure 5: The outline of Daily-Omni Agent workflow.
+3.3 Diagnostic Baseline: Daily-Omni Agent
+To establish a heuristic reference for MLLMs and diagnose the importance of explicit temporal awareness in audio-visual question answering, we constructed a training-free diagnostic pipeline, termed the "Daily-Omni Agent".
+This baseline, designed to understand audio-visual context and answer related questions, integrates several models: Qwen2-Audio [chu2024qwen2audiotechnicalreport], Qwen2.5-VL-7B [bai_qwen25-vl_2025], Whisper-Large-V2 [radford2022robustspeechrecognitionlargescale], and the text-based Qwen2.5-14B-Instruct [qwen2025qwen25technicalreport].
 As shown in Figure 5, when presented with a question, its choices, and the associated video context, the agent first divides both the video and audio streams into three segments of equal duration.
 Subsequently, it independently generates annotations for these segments: visual annotations using Qwen2.5-VL and audio annotations using Qwen2-Audio.
-Additionally, we use Whisper-Large-V2 to provide a transciption of speech and singing of each segment, as Qwen2-Audio tends to omit this information in its annotations.
+Additionally, we use Whisper-Large-V2 to provide a transciption of speech and singing of each segment, as Qwen2-Audio tends to omit this in its annotations.
 Following the segment annotation, the agent utilizes Qwen2.5-VL to perform a consistency check on the visual annotations using the complete video, mirroring the revision process described previously (Section 3.2).
 While generating aligned audio-visual event pairs would ideally provide richer temporal information for reasoning, implementing this step within the agent presents practical challenges.
-Unlike the data curation pipeline, the agent lacks access to a highly capable OLM like Gemini 2.0 Flash for precise event alignment.
-Furthermore, providing a large number of fine-grained event pairs risks overwhelming the context capacity or reasoning capabilities of the agent’s LLM.
-Therefore, as an alternative to full event alignment, we adopt a innovative targeted approach: First, we prompt Qwen2.5-14B-Instruct with the visual and audio annotations, the question, and its choices.
+Unlike the data curation pipeline, the agent lacks access to a highly capable OLM like Gemini 2.5 Flash for precise event alignment.
+Furthermore, providing a large number of fine-grained event pairs may overwhelm the context capacity or reasoning capabilities of the agent’s LLM.
+Therefore, we adopt a targeted approach: First, we prompt Qwen2.5-14B-Instruct with the visual and audio annotations, the question, and its choices.
 The model’s task is to identify a list of specific visual events whose temporal localization is deemed necessary to answer the question correctly.
@@
 4 Experiment
-In this section, we present a comprehensive evaluation of recent open-source and proprietary Multimodal Large Language Models (MLLMs) on our benchmark to delineate their capability boundaries.
+In this section, we present a comprehensive evaluation of recent Multimodal Large Language Models (MLLMs) on our benchmark to delineate their capability boundaries.
 Furthermore, we conduct ablation studies to investigate the key factors influencing model performance.
@@
 Our evaluation encompassed three distinct types of models.
-Firstly, we examined OLMs, including several open-source contenders like VideoLLaMA 2 [cheng_videollama_2024], Unified-IO 2 [lu2023unifiedio2scalingautoregressive], Qwen2.5-Omni [xu_qwen25-omni_2025], Ola [liu_ola_2025] and our Daily-Omni Agent, as well as the proprietary Gemini models [noauthor_introducing_2024].
-Secondly, we assessed VLMs, represented by the open-source Qwen2.5-VL [bai_qwen25-vl_2025] and the proprietary GPT-4o [openai_gpt-4o_2024].
-We also evaluated the performance of some OLMs when provided with only visual inputs.
-Finally, we tested some LLMs such as Deepseek-V3 [deepseekai2025deepseekv3technicalreport], GPT-4o [openai_gpt-4o_2024] and Qwen2.5-14B[qwen2025qwen25technicalreport] without visual and audio inputs to check whether our benchmark contains too much information in questions and choices.
-All model evaluations were conducted strictly according to instructions from their respective official repositories and accompanying documentation (such as "cookbooks" or developer guides).
+Firstly, we examine OLMs, including several open-source contenders like VideoLLaMA 2 [cheng_videollama_2024], Unified-IO 2 [lu2023unifiedio2scalingautoregressive], Qwen2.5-Omni [xu_qwen25-omni_2025], Qwen3-Omni [Qwen3-Omni], Ola [liu_ola_2025] and our Daily-Omni Agent, as well as the proprietary Gemini 2.0 and 2.5 models [noauthor_introducing_2024, gemini25].
+Secondly, we assess VLMs including Qwen2.5-VL [bai_qwen25-vl_2025] and GPT-4o [openai_gpt-4o_2024] and ALMs including Audio Flamingo 3 [goel2025audioflamingo3advancing] and Qwen2-Audio [chu2024qwen2audiotechnicalreport].
+We also evaluate the performance of some OLMs when provided with only visual inputs.
+Finally, we test some LLMs such as Deepseek-V3 [deepseekai2025deepseekv3technicalreport], GPT-4o [openai_gpt-4o_2024] and Qwen2.5-14B[qwen2025qwen25technicalreport] without visual and audio inputs to evaluate the extent of textual bias in our questions and choices.
+All model evaluations are conducted strictly according to instructions from their respective official repositories and accompanying documentation (such as "cookbooks" or developer guides).
 4.2 Main Results
-Table 2 presents comprehensive evaluation results, shedding light on the real-world audio-visual understanding capabilities of contemporary MLLMs and our proposed Daily-Omni Agent.
-Firstly, earlier Omni-modal Language Models (OLMs) such as Unified-IO 2 and VideoLLaMA 2 demonstrate limited performance on our benchmarks.
+Table 2: Performance comparison of MLLMs on Daily-Omni.
+Boldface and underline indicate the top two performers.
+Subscripts on ’Avg’ for visual-only OLMs show the performance drop from their audio-visual counterparts.
+Abbreviations: AV Align = audio-visual alignment, Comp.
+= comparative, Ctx.
+Und.
+= context understanding, Evt.
+Seq.
+= event sequence, Infer.
+= inference, Reas.
+= reasoning, and 30s/60s denote duration-based subsets.
+Random guess accuracy is 25%.
+Color code: Gray background indicates closed-source models and blue background indicates open-source models.
+Methods AV Align Comp. Ctx. Und. Evt. Seq. Infer. Reas. 30s 60s Avg
+Omni-Modal Language Models (With Visual and Audio)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Omni-Modal Language Models (Visual Only)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Omni-Modal Language Models (Audio Only)
+[…]
+[…]
+[…]
+[…]
+Visual Language Models (Visual Only)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Audio Language Models (Audio Only)
+[…]
+[…]
+Textual Language Models (Without Visual and Audio)
+[…]
+[…]
+[…]
+Table 2 presents comprehensive evaluation results, shedding light on the real-world audio-visual understanding capabilities of contemporary MLLMs and our diagnostic reference baseline, the Daily-Omni Agent.
+Firstly, earlier Omni-modal Language Models (OLMs) such as Unified-IO 2 and VideoLLaMA 2 demonstrate limited performance on our benchmark.
 Notably, their results are, in several instances, even surpassed by text-only LLMs.
 Furthermore, the Unified-IO 2 series displays a perplexing degradation in performance with increasing model size, an observation consistent with findings from OmniBench [li_omnibench_2024].
-A plausible explanation is an inadequate capacity within these language models to effectively process and synthesize cross-modal information.
-Secondly, while overall performance still presents challenges, more recent open-source OLMs such as the Qwen2.5-Omni series and Ola demonstrate reasonable proficiency in inference and reasoning tasks.
-This suggests a foundational understanding of general visual and audio contexts.
-However, their efficacy significantly diminishes on temporally-sensitive tasks like audio-visual event alignment and context understanding, where results remain largely unsatisfactory.
-Despite incorporating cross-modal positional embeddings designed to provide temporal awareness, these models still struggle to address questions requiring precise temporal capabilities.
-Conversely, proprietary OLMs such as Gemini 2.0 Flash exhibit markedly superior cross-modal temporal capabilities, attaining the leading overall performance score of 67.84%.
-The observation that even top-tier OLMs have not yet breached the 70% accuracy threshold underscores the demanding nature of this benchmark, positing it as a valuable and pertinent objective for the advancement of contemporary OLMs.
-Thirdly, upon ablating audio input and providing only visual data, all evaluated OLMs exhibit a substantial decline in performance.
-Notably, the magnitude of this performance loss tends to be greater for models that demonstrated superior initial performance when leveraging both modalities.
-The significant performance drop, especially in high-performing models, when deprived of audio, validates that the tasks within this benchmark genuinely require and benefit from the integration of both auditory and visual information, rather than being solvable by visual cues alone.
-This dependence on both modalities is further evidenced by the performance of leading Visual Language Models (VLMs).
-For instance, Qwen2.5-VL (7B) achieves only 40.68% accuracy, and even the highly capable GPT-4o (with its visual input) reaches just 56.47%.
-Given that these are state-of-the-art models within their respective size categories (sub-7B and larger proprietary models), their comparatively modest performance on our benchmark further underscores its value in demanding genuine audio-visual integration, a capability absent in previous multi-modal benchmarks.
-These results reveal that: (1) Current OLMs demonstrate reasonable general audio and visual understanding capabilities when processing videos under 60 seconds in duration.
-However, they continue to struggle with tasks demanding sophisticated cross-modal temporal awareness or the integration of information across extended time intervals.
-(2) Our proposed question-answer generation pipeline proves effective in creating challenging evaluation instances that accurately probe the audio-visual understanding capabilities of MLLMs.
-Table 2: Performance comparison of various MLLMs across different question domains and video durations.
-For Omni-modal Language Models (OLMs) with audio and visual inputs, Boldface shows the best performance and underlines show the second best.
-For OLMs with visual inputs only, subscripts on the ’Avg’ score demonstrate the performance difference compared to their audio-visual counterparts.
-For your reference, the accuracy rate of pure guessing is 25%.
-More details of model evaluation are provided in Appendix C.
-Methods AV Event Alignment Comparative Context Understanding Event Sequence Inference Reasoning 30s Subset 60s Subset Avg
-Omni-Modal Language Models (With Visual and Audio)
-Qwen2.5-Omni (7B) $44.12$ $51.15$ $38.86$ $40.52$ $57.79$ $61.71$ $46.68$ $48.36$ $47.45$
-Qwen2.5-Omni (3B) $38.66$ $48.09$ $33.68$ $33.99$ $54.55$ $44.00$ $42.35$ $38.36$ $40.52$
-Ola (7B) $40.34$ $61.07$ $40.41$ $43.46$ $63.64$ $69.71$ $51.47$ $49.82$ $50.71$
-Unified-IO-2 L (1B) $27.31$ $22.90$ $26.42$ $27.78$ $29.87$ $29.14$ $27.67$ $27.09$ $27.40$
-Unified-IO-2 XL (3B) $30.25$ $30.53$ $25.39$ $29.08$ $33.12$ $21.71$ $28.13$ $28.55$ $28.32$
-Unified-IO-2 XXL (8B) $25.63$ $31.30$ $26.42$ $25.82$ $35.06$ $29.71$ $26.74$ $30.00$ $28.24$
-VideoLLaMA2 (7B) $35.71$ $35.88$ $35.75$ $31.70$ $40.91$ $34.29$ $38.02$ $31.82$ $35.17$
-Gemini 2.0 Flash 62.18 73.28 63.73 63.72 76.62 75.43 67.23 68.55 67.84
-Gemini 2.0 Flash Lite 55.04 $64.89$ $58.03$ 54.25 $74.03$ 72.00 $62.44$ 60.00 $61.32$
-Daily-Omni (ours) $51.68$ 68.70 60.10 $53.92$ 78.57 $71.43$ 63.99 $59.27$ 61.82
-Omni-Modal Language Models (Visual Only)
-Qwen2.5-Omni (7B) $38.24$ $48.85$ $34.72$ $36.27$ $51.95$ $45.71$ $40.80$ $41.64$ $41.19_{\text{-}6.3}$
-Qwen2.5-Omni (3B) $33.61$ $42.75$ $36.27$ $33.33$ $49.35$ $38.86$ $38.79$ $36.55$ $37.76_{\text{-}2.8}$
-Gemini 2.0 Flash 39.08 64.12 56.48 56.21 67.53 $62.29$ $56.57$ $55.45$ 56.06-${}_{\text{-}11.8}$
-Gemini 2.0 Flash Lite $43.70$ $58.02$ $53.89$ $45.10$ $64.29$ $60.57$ $53.01$ $51.64$ 52.38-${}_{\text{-}8.9}$
-Visual Language Models (Visual Only)
-GPT-4o $47.90$ $62.60$ $52.33$ $52.61$ 66.23 66.29 55.64 57.45 56.47
-Qwen2.5-VL (7B) 36.97 46.56 $33.68$ 37.91 51.95 44.00 39.26 42.36 40.68
-Qwen2.5-VL (3B) $35.71$ $43.51$ 34.72 $33.66$ $43.51$ $39.43$ $37.71$ $37.09$ $37.43$
-Textual Language Models (Without Visual and Audio)
-GPT-4o $33.19$ $43.51$ $28.50$ $30.39$ $44.81$ $46.86$ $36.48$ $36.18$ $36.34$
-Deepseek-V3 (671B) $31.93$ $41.22$ $29.02$ $29.41$ $44.81$ $46.29$ $35.24$ $36.00$ $35.59$
-Qwen2.5-Instruct (14B) $30.25$ $39.69$ $27.98$ $28.43$ $42.21$ $42.86$ $32.15$ $35.82$ $33.83$
-4.3 Results of Daily-Omni Agent
-Our Daily-Omni Agent, leveraging Qwen2.5-VL-7B, Qwen2-Audio (7B), and Qwen2.5-Instruct-14B, achieves 61.82% overall accuracy.
-This performance is state-of-the-art among open-source methods and surpasses smaller proprietary MLLMs.
-The success of our agent suggests that by simply leveraging VLM and ALM with time-period-level alignment coupled with event-level alignment for some key events, we are able to create a very strong OLM.
+A plausible explanation is that these models may not reliably synthesize cross-modal information under our evaluation setting.
+Secondly, we observe a clear performance divide among recent OLMs, highlighting the importance of cross-modal temporal alignment for audio-visual QA.
+Our training-free Daily-Omni Agent, which makes temporally localized evidence more explicit via a decoupled pipeline, serves as a useful reference point with an overall accuracy of 61.82%.
+Models that perform substantially below this reference (such as Ola and earlier OLMs) appear to struggle with synchronizing parallel audio and visual streams for alignment-sensitive questions.
+In contrast, models that achieve higher performance demonstrate stronger proficiency in audio-visual reasoning.
+While proprietary models like Gemini 2.5 Flash achieve top-tier performance (up to 73.06%), among open-source models, the Qwen2.5-Omni and Qwen3-Omni series stand out.
+Their gains likely reflect a combination of factors, including training data and objectives, architectural choices for multimodal fusion, and temporal modeling components.
+In particular, explicit cross-modal positional designs (e.g., TMRoPE) may be one contributing technique for improving temporal awareness across modalities.
+Overall, these results suggest that robust spatio-temporal alignment mechanisms remain an important ingredient for tackling complex audio-visual reasoning tasks.
+These results reveal that: (1) Advanced OLMs with stronger temporal handling (e.g., the Qwen Omni series) or proprietary systems (e.g., Gemini) demonstrate solid audio-visual understanding, though they still face challenges in complex temporal reasoning.
+(2) Models with weaker temporal integration tend to perform poorly on alignment-sensitive questions, even when provided with both audio and visual inputs.
+(3) Our QA construction and filtering pipeline produces challenging evaluation instances that probe the temporal reasoning boundaries of current MLLMs.
+Figure 6: Evaluation stability under video-level subsampling. Lines show median overall accuracy and shaded bands denote the 5th–95th percentile across 200 trials.
+4.3 Evaluation Stability via Subsampling
+To assess the stability of aggregate evaluation on Daily-Omni, we conduct a video-level subsampling study on representative OLMs.
+We repeatedly sample {20,40,60,80}% of videos without replacement and evaluate overall accuracy on the induced QA subsets (Figure 6).
+The estimates remain consistent across subsamples and the variability decreases rapidly as more videos are included: at 80% of videos (approximately 958 questions), the 5th–95th percentile range is within 1.1–1.2 percentage points for Gemini 2.5 Flash, Gemini 2.5 Flash Lite, and Qwen3-Omni-30B.
+Overall, this subsampling analysis supports that Daily-Omni provides sufficiently stable aggregate evaluation to serve as a robust benchmark.
+4.4 Ablation Study
+To further validate the design of our proposed baseline and the intrinsic multimodal requirements of the Daily-Omni benchmark, we conduct comprehensive ablation studies focusing on temporal alignment methods and modality dependency.
+(a) Modality ablation on Daily-Omni
+(b) Impact of alignment methods
+Figure 7: Ablation studies validating the design of the benchmark and the baseline agent.
+(a) Performance comparison of representative OLMs under full modalities (A+V), Audio-only, and Visual-only settings.
+(b) Comparison of average accuracy (bars) and aligned event pairs per question (line) using our baseline agent.
+4.4.1 Modality Ablation
+To verify whether the tasks in Daily-Omni genuinely require the integration of both auditory and visual information, we evaluated representative top-performing OLMs (Qwen3-Omni-30B-A3B-Instruct/Thinking, Gemini 2.5 Flash, and Gemini 2.5 Flash Lite) under three conditions: full modalities (Audio + Visual), audio-only, and visual-only.
+The comparative results are illustrated in Figure 7(a).
+Upon ablating either modality, all evaluated OLMs exhibit a substantial decline in performance.
+For instance, the highly capable Gemini 2.5 Flash experiences a drastic accuracy drop from 73.06% to 54.05% without visual input, and down to 44.61% without audio input.
+Similarly, the Qwen3-Omni models suffer a 13–16% performance degradation when deprived of either modality.
+This significant performance gap between the full-modality and uni-modal settings strongly validates that the questions in our benchmark cannot be reliably answered using a single sensory channel or text-based guessing.
+Interestingly, we observe that for these models, the audio-only performance is generally higher than the visual-only performance on our dataset.
+This observation highlights that unlike many existing video QA benchmarks that exhibit a strong visual bias (where audio is often redundant), Daily-Omni contains rich, indispensable acoustic events and spoken dialogues.
+In conclusion, this ablation confirms that achieving high performance on Daily-Omni strictly necessitates genuine multi-modal synergy and precise cross-modal integration.
+4.4.2 Daily-Omni Agent Alignment
+Our heuristic Daily-Omni Agent, leveraging Qwen2.5-VL-7B, Qwen2-Audio (7B), and Qwen2.5-Instruct-14B, achieves 61.82% overall accuracy.
+The fact that this simple, decoupled pipeline surpasses smaller proprietary MLLMs and several unified open-source models strongly validates the core motivation of our benchmark: temporal alignment is the primary bottleneck in audio-visual reasoning.
 To further study the effect of event alignment, we conducted an ablation study evaluating the Daily-Omni Agent’s performance under three alignment scenarios: (1) No Alignment: Generating comprehensive visual and audio captions as a base, but omitting any form of event alignment.
-(2) Naive Alignment: Generating comprehensive visual and audio captions, followed by a step where the VLM is prompted to extract question-relevant events and their temporal boundaries (begin/end times).
-The ALM is then utilized to produce audio captions specifically for those extracted segments and attaining aligned audio-visual event pairs in the process.
+(2) Naive Alignment: Generating comprehensive visual and audio captions, then aligning events in a question-agnostic manner.
+Specifically, we prompt the VLM to enumerate visual events from the captions and localize each event by predicting its temporal boundaries (start/end).
+For each localized interval, we retrieve the corresponding audio events within the same time range using the ALM, and form aligned audio–visual event pairs.
 (3) Smart Alignment: Generating comprehensive visual and audio captions and aligned events as stated in Section 3.3.
 Note that for (2) and (3), we only consider generating an align event pair when the provided segment has a duration below a certain threshold to prevent confusion.
-The overall accuracy and the average number of aligned event pairs identified per question for each method are presented in Table 3.
+The overall accuracy and the average number of aligned event pairs identified per question for each method are presented in Figure 7(b).
 As expected, the Smart Alignment method achieves the highest average accuracy.
@@
 Consequently, we hypothesize that equipping the agent with a more powerful open-vocabulary video temporal grounding model would unlock further significant improvements in its performance, mitigating the impact of imprecise temporal grounding.
-Figure 6: MLLMs’ accuracy over different question categories
-Align Method Average Accuracy Aligned Event Pairs per Question
-No Alignment $60.65$ 0
-Naive Alignment $59.65$ 0.68
-Smart Alignment 61.82 1.11
-Table 3: Impact of different aligning methods on average accuracy and aligned event pairs per question.
-4.4 Case Study
-This part delves into selected case studies of model performance on Daily-Omni.
-Such an analysis of specific examples offers deeper insights into both the nature of the Daily-Omni benchmark and the current capabilities and limitations of the models.
-Figure 7 presents three illustrative questions from the Daily-Omni benchmark.
-It displays the responses of the Gemini 2.0 Flash and Qwen2.5-Omni models under two conditions: processing full audio-visual inputs versus visual-only inputs.
-Figure 7: OLM responses to Daily-Omni questions: A comparison of audio-visual versus visual-only inputs.
-The figure presents examples of OLM performance when using full audio-visual modalities compared to visual-only input.
-For each case, it shows the audio-visual content, the question with choices (correct answer highlighted), and the models’ answers, indicating correctness for each input condition.
-Case 1: Chronological Event Ordering This question requires determining the first event to occur in the video from a given list of audio-visual events.
-To correctly ascertain the existence and chronological position of choices B (Player discussing a personal milestone) and D (Female presenter announcing a groundbreaking sports contract), the model must leverage both audio and visual sensory inputs.
-Gemini 2.0 Flash successfully identifies choice B as the initial event, whereas Qwen2.5-Omni 7B fails to do so.
-Both model failed to answer correctly when only taking visual input.
-This disparity suggests that merely possessing multi-modal sensory capabilities is insufficient; models also require robust cross-modal reasoning to effectively integrate audio-visual information and achieve correct temporal ordering.
-Case 2: Audio-Visual Inference This case assesses the ability to discern the functional differences between the hand gestures of two individuals engaged in a discussion.
-While visual input captures the gestures themselves, audio input is also critical for interpreting their communicative purpose as it reveals what is being said and who said it.
-Gemini 2.0 Flash correctly identified the functional difference (D) with audio-visual input but failed with visual-only input, selecting a more generic but incorrect interpretation (A).
-Qwen2.5-Omni 7B, failed to make the correct distinction in both audio-visual and visual-only conditions, selecting an unreasonable option (B).
-This question underscores that merely understanding the speech content is insufficient; accurately identifying the speaker for each utterance is also crucial.
-Successfully attributing speech requires robust audio-visual temporal integration—the ability to precisely synchronize the spoken audio with the visual depiction of the active speaker.
-Following this, strong reasoning ability is needed to infer the distinct roles and communicative intents (e.g., inquirer vs.
-narrator) based on the combined audio-visual information, and thereby understand the different functions of their gestures.
-The failure of Qwen2.5-Omni 7B, even with audio-visual input, suggests potential limitations in one or both of these sophisticated cross-modal capabilities.
-Case 3: Sound Understanding This question directly tests the model’s ability to analyze and compare audio characteristics across two different segments of a video-an skateboard demonstration and an outdoor skatepark scene.
-The task involves identifying distinct audio elements such as human speech, specific sound effects (e.g., skateboard impact), and background music.
-This is fundamentally an audio-centric task where visual context primarily helps situate the sounds.
-Strikingly, both Gemini 2.0 Flash and Qwen2.5-Omni failed to correctly compare the audio content, selecting the same incorrect option (D) irrespective of whether they received audio-visual or visual-only input.
-Their failure in the audio-visual condition suggests a significant challenge in fine-grained auditory scene analysis, such as distinguishing non-verbal sound event and music, or accurately mapping these perceived audio features to the descriptive choices provided.
 5 Conclusion
-This paper introduced Daily-Omni, a novel Audio-Visual Question Answering benchmark designed to evaluate MLLMs on temporally-aligned multimodal reasoning in daily life scenarios.
-We also proposed an efficient Daily-Omni QA Generation Pipeline and the training-free Daily-Omni Agent, a strong open-source baseline.
-Our evaluation revealed that while recent MLLMs show general audio-visual understanding, they significantly struggle with precise cross-modal temporal awareness The results underscore our pipeline’s value and effectiveness in creating such challenging questions.
-The Daily-Omni Agent, with its targeted temporal alignment, achieved competitive results, and our ablation study confirmed that a Smart Alignment strategy improves performance, unlike naive approaches.
-In summary, Daily-Omni highlights the current MLLM frontiers in complex audio-visual temporal reasoning.
-Future efforts should prioritize developing more accurate and robust multimodal temporal grounding techniques for truly sophisticated audio-visual understanding.
-Appendix A Societal Impact
-The Daily-Omni benchmark and the associated research aim to advance the capabilities of Multimodal Large Language Models (MLLMs) in understanding and reasoning about real-world audio-visual scenarios.
-Positive Societal Impacts:
-  • Improved Human-AI Interaction: MLLMs with better audio-visual understanding can lead to more natural and effective human-AI interaction in various applications, such as assistive technologies for individuals with disabilities (e.g., describing surroundings for visually impaired individuals), smarter virtual assistants, and more intuitive robotics.
-  • Enhanced Content Accessibility and Analysis: Advanced MLLMs can automatically generate rich descriptions, summaries, and answer questions about video content, making vast amounts of audio-visual data more accessible and searchable.
-This can benefit education, media archiving, and content moderation.
-  • Development of Artificial Agent Systems capable of perceiving and understanding the outcome of their decision or operation via audio and visual modalities will be smarter.
-Negative Societal Impacts and Mitigation:
-  • Surveillance and Privacy Concerns: Enhanced audio-visual understanding capabilities could be deployed in surveillance systems, raising significant privacy concerns if not implemented with strong ethical guidelines and legal oversight.
-Mitigation: Development and deployment of such technologies must be accompanied by robust privacy-preserving techniques (e.g., federated learning, on-device processing, data anonymization) and strong regulatory frameworks that protect individual privacy.
-The videos in Daily-Omni are sourced from publicly available platforms, but the broader application of the technology requires careful consideration.
-  • Bias Amplification: MLLMs are trained on large datasets, which may contain societal biases.
-These biases can be learned and amplified by the models, leading to unfair or discriminatory outcomes in their applications.
-Mitigation: Efforts in dataset curation for Daily-Omni aimed for diversity in video categories (Section 3.1, Figure 2).
-Ongoing research into bias detection, fairness metrics, and debiasing techniques for MLLMs is essential.
-Continuous auditing of model performance across different demographic groups is also necessary.
-Our work primarily focuses on creating a benchmark to evaluate and drive progress in MLLM capabilities.
-We believe that by transparently assessing these models, we can better understand their strengths and weaknesses, which is a prerequisite for responsible development and deployment.
-Appendix B Licenses for Existing Assets
-This research utilizes several existing assets, including datasets and pre-trained models.
-We have made efforts to properly credit the original creators through citations in the main paper and adhere to their terms of use.
-Datasets Used for Video Sourcing:
-  • AudioSet [Audioset]: The AudioSet dataset consists of an expanding ontology of 632 audio event classes and a collection of 2,084,320 human-labeled 10-second sound clips drawn from YouTube videos.
-The video segments used in Daily-Omni were derived from the original full-length YouTube videos corresponding to AudioSet clips.
-AudioSet itself is released by Google under a Creative Commons license (typically CC BY 4.0 for the annotations, while the YouTube videos are subject to their own licenses set by uploaders).
-We accessed publicly available YouTube videos.
-  • Video-MME [fu2024videommefirstevercomprehensiveevaluation]: Video-MME is a benchmark for evaluating multimodal models on video understanding.
-The videos are sourced from YouTube.
-The benchmark is only used for academic research.
-Commercial use in any form is prohibited.
-  • FineVideo [Farré2024FineVideo]: FineVideo is a dataset for fine-grained video content analysis.
-The videos provided are derived from YouTube-Commons.
-All the videos are shared under a CC-By license and, in accordance with that license, every YouTube channel is fully credited.
-For all video data sourced from YouTube, we rely on the videos being publicly accessible and adhere to YouTube’s Terms of Service.
-The specific license for each YouTube video is determined by its uploader.
-Our use is for non-commercial research purposes to create a benchmark.
-Models Used (Annotation, QA Generation, Daily-Omni Agent, Evaluation):
-  • Whisper [radford2022robustspeechrecognitionlargescale]: Developed by OpenAI.
-The Whisper models are typically released under a permissive license (e.g., MIT License), allowing for research and commercial use.
-We used Whisper-Large-V2 and Whisper-V3-Large.
-  • Gemini Models [noauthor_introducing_2024]: Developed by Google.
-Access to Gemini models (e.g., Gemini 2.0 Flash, Gemini 2.0 Flash Lite) is typically through Google’s APIs (e.g., Google AI Studio or Vertex AI), subject to Google’s terms of service and API usage policies.
-  • Deepseek Models (Deepseek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability], Deepseek-V3 [deepseekai2025deepseekv3technicalreport]): Developed by DeepSeek AI.
-These models are often available via API or with specific research/community licenses.
-Their terms of use were respected.
-  • Qwen Models (Qwen2-Audio [chu2024qwen2audiotechnicalreport], Qwen2.5-VL [bai_qwen25-vl_2025], Qwen2.5-Omni [xu_qwen25-omni_2025], Qwen2.5-14B-Instruct [qwen2025qwen25technicalreport]): Developed by Alibaba Cloud.
-Qwen models are often released under licenses like the Tongyi Qianwen LICENSE AGREEMENT, which may have specific restrictions for commercial use but generally permit research.
-  • GPT-4o [openai_gpt-4o_2024]: Developed by OpenAI. Accessed via OpenAI’s API, subject to their terms of service and usage policies.
-  • Other MLLMs for Evaluation (VideoLLaMA 2 [cheng_videollama_2024], Unified-IO 2 [lu2023unifiedio2scalingautoregressive], Ola [liu_ola_2025]): These models are from various research institutions and companies.
-Their use in our evaluation adheres to the licenses provided with their respective releases (e.g., Apache 2.0, MIT, or custom research licenses).
-Original papers are cited.
-We have strived to use all assets in accordance with their intended purpose for research and benchmarking, and have credited original authors.
-The Daily-Omni benchmark itself, including the QA pairs, will be released under CC BY-NC-SA 4.0 to encourage its use and further development in the research community, while respecting the licenses of the underlying video content.
-Appendix C Evaluation Details
-The models listed in Table 2 were evaluated as follows: Qwen2.5-Omni (7B and 3B), Ola, Unified-IO-2 (L, XL, and XXL), VideoLLaMA2, and Qwen2.5-VL (7B and 3B) were deployed and tested on a local server.
-In contrast, Gemini 2.0 Flash, Gemini 2.0 Flash Lite, GPT-4o, and Qwen2.5-Instruct (14B) were evaluated via their respective APIs.
-For the Daily-Omni Agent, Qwen2.5-VL-7B-Instruct and Qwen2.5-14B-Instruct were accessed via the API provided by Alibaba Cloud’s Bailian platform (accessible at https://bailian.console.aliyun.com/) to ensure operational efficiency.
-Similarly, Whisper-Large-V2 was utilized via the API provided by OpenAI.
-The Qwen2-Audio component, on the other hand, was deployed and tested locally.
-It is important to note that all component models of Daily-Omni are capable of local deployment.
-According to official documentation from Aliyun (https://help.aliyun.com/zh/model-studio/model-user-guide/) and OpenAI (https://openai.com/index/introducing-chatgpt-and-whisper-apis/), the Qwen models offered through Bailian and the Whisper-Large-V2 model accessed via the OpenAI API are identical to their respective open-source or standard versions.
-Therefore, no performance differences are anticipated.
-Our code implements direct passing of local video path to the Qwen2.5-VL API.
-However, this functionality might require you to contact Aliyun customer service to enable.
-If direct path input is not activated, you can alternatively pass a list of video frames, though this may result in suboptimal performance.
-[…]
+This paper introduced Daily-Omni, an Audio-Visual Question Answering benchmark for evaluating temporally aligned multimodal reasoning in daily-life scenarios, together with a scalable QA construction and verification pipeline.
+Our evaluation of 24 foundation models under 37 model–modality settings shows that while recent MLLMs exhibit strong unimodal capabilities, they still struggle on questions that require fine-grained cross-modal temporal alignment.
+We also include a training-free modular reference baseline (Daily-Omni Agent) to help interpret alignment-sensitive failure modes and to illustrate the impact of making temporally localized evidence explicit.
+Overall, Daily-Omni highlights temporal alignment as a key bottleneck for current unified architectures.
+Future work should prioritize more accurate and robust multimodal temporal grounding and alignment mechanisms for audio-visual reasoning.